### v3

# A Robust Mixed-Effects Bandit Algorithm for Assessing Mobile Health Interventions

Abstract
Mobile health leverages personalized, contextually-tailored interventions optimized through bandit and reinforcement learning algorithms. Despite its promise, challenges like participant heterogeneity, nonstationarity, and nonlinearity in rewards hinder algorithm performance. We propose a robust contextual bandit algorithm, termed “DML-TS-NNR”, that simultaneously addresses these challenges via (1) modeling the differential reward with user- and time-specific incidental parameters, (2) network cohesion penalties, and (3) debiased machine learning for flexible estimation of baseline rewards. We establish a high-probability regret bound that depends solely on the dimension of the differential reward model. This feature enables us to achieve robust regret bounds even when the baseline reward is highly complex. We demonstrate the superior performance of the DML-TS-NNR algorithm in a simulation and two off-policy evaluation studies.
1 Introduction
Mobile health (mHealth) uses smart devices to deliver digital notification interventions to users. These notifications nudge users toward healthier attitudes and behaviors. Because mHealth applications can monitor and react to users and their environment, they offer the promise of personalized, contextually-tailored interventions. In practice, achieving this promise requires bandit or reinforcement learning (RL) algorithms that can accurately learn mHealth intervention effects, including how effects vary by user, over time, and based on context. In this regard, contextual bandit algorithms are appealing due to strong empirical performance in other settings, their ability to customize intervention decisions based on changing contexts, and their simplicity and extensibility relative to full RL algorithms.
One unique aspect of mHealth (Greenewald et al., 2017) is the presence of a “do-nothing” action in which the mHealth app does not intervene. As mHealth rewards typically involve human decision making (e.g., whether the user chooses to exercise), the distribution of the the corresponding baseline reward—the reward under the “do-nothing” action—typically evolves in a complex, nonstationary fashion as time progresses and contexts change. In contrast, the treatment effects—the difference in the expected value of the reward relative to the “do-nothing” action—tend to be much more stable and can be adequately modeled using classical stationary models; e.g., linear models (Robinson, 1988).
Greenewald et al. (2017) introduced an action-centered contextual bandit algorithm with sub-linear regret in this setting by replacing the observed reward, $R_{t}$, with a “pseudo-reward,” $\tilde{R}_{t}$, where $t$ indexes the time points. Denoting the binary action as $A_{t}\in\{0,1\}$ with $A_{t}=0$ corresponding to the “do-nothing” action, the pseudo-reward for $A_{t}$ can be written as
$\tilde{R}_{t}:=(A_{t}-\pi_{t})R_{t}=\pi_{t}(1-\pi_{t})R_{t}\left(\frac{A_{t}}{\pi_{t}}-\frac{1-A_{t}}{1-\pi_{t}}\right),$
where $\pi_{t}$ is the (known) probability of treatment assignment from the Thompson-sampling algorithm. We took an additional step to derive an equivalent expression (second equality), which reveals that $\tilde{R}_{t}$ is proportional to an inverse-probability-weighted (IPW) estimator. Crucially, $\tilde{R}_{t}$ provides an unbiased estimate of the treatment effect regardless of the non-stationarity of the baseline outcome.
Although Greenewald et al. (2017)’s analysis shows sub-linear regret in nonstationary settings, our simulations show that it is outperformed by other methods over finite time horizons. The reasons include its failure to address treatment effect heterogeneity, pool information across users, and model the baseline outcome, which leads to highly variable pseudo-rewards and slow learning. We propose to generalize their method to overcome these limitations by
  1. 1.
    imposing hierarchical structure in the treatment effect model with shared structural parameters and “incidental parameters” for users and time,
  2. 2.
    efficiently pooling across users and time via nearest-neighbor regularization (NNR),
  3. 3.
    de-noising rewards with flexible supervised learning models via the de-biased machine learning (DML) framework of Chernozhukov et al. (2018).
We name the resulting method DML-TS-NNR to highlight its usage of DML, NNR, and Thompson Sampling. The primary contributions are (a) the DML-TS-NNR method, (b) a high-probability regret bound that relies solely on the dimension of the differential reward model, which is typically much smaller than that of the complex baseline reward model, and (c) empirical comparisons demonstrating the superior performance of DML-TS-NNR in simulation and two real-world mHealth studies.
2 Related Work
Related work in statistics considers treatment effect estimation with longitudinal data. Cho et al. (2017) present a semiparametric random effects method for estimating subject-specific static treatment effects. Qian et al. (2020) discusses the statistical challenges of applying linear mixed-effects models to mHealth data and shows that the resulting estimates may lack a causal interpretation.
In the bandit literature, several works address a growing pool of users. Ma et al. (2011) introduced NNR in recommender systems using homophily principles—similar nodes are more likely to be connected than dissimilar ones (McPherson et al., 2001). Cesa-Bianchi et al. (2013) adapted NNR for bandit settings, improving regret. Subsequent improvements focused on scalability and algorithm modifications for stronger regret bounds (Vaswani et al., 2017; Yang et al., 2020).
Other bandit approaches explicitly address the longitudinal setting in which treatment effects may evolve over time. The intelligentpooling method of Tomkins et al. (2021) employs a Gaussian linear mixed-effects model with both user- and time-specific random effects. Hu et al. (2021) provides a related approach based on generalized linear mixed effects models. Unlike our approach, both methods require the (generalized) linear conditional mean model to be correctly specified.
Aouali et al. (2023) and Lee et al. (2024) also propose bandit algorithms with mixed effects; methodologically, however, their use of mixed effects is dissimilar from that of the works above and our own. Aouali et al. (2023) employs random effects to relate treatments with categories (e.g., movies within a genre), and Lee et al. (2024) considers the case in which a single, discrete action leads to a multivariate reward that is affected by a user-specific random effect. In contrast, the other approaches listed above (and our own) do not assume any relationships among the actions, and they associate each discrete action with a single scalar-valued reward in an iterative fashion.
Other work develops semiparametric methods that lead to sublinear regret without requiring correct specification of the conditional mean reward. Krishnamurthy et al. (2018) and Kim & Paik (2019) propose improvements to the approach of Greenewald et al. (2017) that lead to stronger regret bounds. Kim et al. (2021, 2023) develop a doubly-robust approach for both linear and generalized linear contextual bandits. However, in contrast to Tomkins et al. (2021), Hu et al. (2021), and our approach, these semiparametric methods are not immediately applicable to longitudinal settings.
3 Setting
3.1 Problem Statement
Recognizing the connection between mHealth policy learning and contextual bandit methods, we now consider a contextual multi-armed bandit environment with one control arm ($a=0$) and $q$ non-baseline arms corresponding to different actions or treatments. Individuals are indexed by $i=1,2,\ldots$, and decision points are indexed by $t=1,2,\ldots$. For each individual $i$ at time $t$, the algorithm observes a context vector $S_{i,t}\in\mathcal{S}$, chooses an action $A_{i,t}\in[q]:=\{0,\ldots,q\}$, and receives a reward $R_{i,t}\in\mathbb{R}$. The differential reward describes the difference in expected rewards between choosing a non-baseline arm and the control arm, defined as $\Delta_{i,t}(s,a):=\mathbb{E}\left[R_{i,t}|S_{i,t}=s,A_{i,t}=a\right]-\mathbb{E}\left[R_{i,t}|S_{i,t}=s,A_{i,t}=0\right]$.
We denote the observation history up to time $t$ for individual $i$ by $\mathcal{H}_{i,t}:=(S_{i1},A_{i1},R_{i1},\dots,R_{it})$. Actions are selected according to stochastic policies, $\pi_{i,t}:\mathcal{H}_{i,t}\times\mathcal{S}\to\mathcal{P}([q])$. We use $\pi_{i,t}(a|s)$ to denote the probability of action $a\in[q]$ given current context $s\in\mathcal{S}$ for a fixed (implicit) history. As in Greenewald et al. (2017), we assume that the probability of control action is bounded:
Assumption 1.
There exists $10<\pi_{\min},\pi_{\max}<$ 1 s.t $\pi_{\min}<\pi_{i,t}(0|a),1-\pi_{i,t}(0|a)<\pi_{\max}\forall i,t,a$.
As Greenewald et al. (2017) noted, this assignment probability controls the number of messages that participants receive. It ensures that participants receive sufficient messages that they do not disengage, but not so many that they are overwhelmed or fatigued. We now define
$A^{*}=\operatorname{argmax}_{a\in[q]}\mathbb{E}\left[R_{i,t}|S_{i,t},A_{i,t}=a\right],\quad\bar{A}^{*}=\operatorname{argmax}_{a\in[q]\backslash\{0\}}\Delta_{i,t}(S_{i,t},a),$
the optimal treatment arm (including control) and the alternative arm, respectively. Respecting the restrictions in Assumption 1, we can now express the optimal policy, $\pi_{i,t}^{*}$, in two cases:
  1. 1.
    $A^{*}=0$: $\pi_{i,t}^{*}(0|S_{i,t})=\pi_{\max}$ and $\pi_{i,t}^{*}(\bar{A}^{*}|S_{i,t})=1-\pi_{\max}$.
  2. 2.
    $A^{*}>0$: $\pi_{i,t}^{*}(A^{*}|S_{i,t})=1-\pi_{\min}$ and $\pi_{i,t}^{*}(0|S_{i,t})=\pi_{\min}$.
These cases represent achieving the appropriate boundary condition given the optimal action, i.e., maximum chance of “do-nothing” action if $A^{\star}=0$ and minimum chance if $A^{\star}\not=0$. Otherwise, $\pi_{i,t}^{*}(a|S_{i,t})=0$. In Section 5.2, we define regret relative to this optimal policy, $\pi_{i,t}^{*}$.
Lastly, we consider a study design that progresses in stages, where individuals enter the study sequentially (Friedman et al., 2010), as illustrated in Figure 3 in Appendix A. Staged recruitment serves two purposes in our analysis: it increases the fidelity of the theoretical setting by mimicking real-world recruitment strategies, and it allows us to estimate changes in the differential reward over time. Concretely, we first observe individual $i=1$ at time $t=1$. Then we observe individuals $i=(1,2)$ at times $t=(2,1)$. Subsequent stages $k$ involve individuals $i\leq k$, each having had $k-i+1$ observations, respectively. Define $\mathcal{O}_{k}=\{(i,t):i\leq k\>\&\>t\leq k+1-i\}$ as the double index set of all observations in stage $k$. A notation summary is presented in Appendix G.
3.2 Doubly Robust Differential Reward
Following Greenewald et al. (2017), we assume that the differential reward is linear:
$\Delta_{i,t}(s,a)=x(s,a)^{\top}\theta_{i,t},$ (1)
where $x(s,a)\in\mathbb{R}^{p\times 1}$ is a feature vector depending on the context and action. We allow the baseline reward, $g_{t}(s)$, to be a complex, nonlinear, and nonstationary function of context and time. Combined, these two assumptions imply a partially linear model for the conditional mean reward:
$\begin{split}\mathbb{E}\left[R_{i,t}|S_{i,t}=s,A_{i,t}=a\right]&=x(s,a)^{\top}\theta_{i,t}\delta_{a>0}+g_{t}(s),\end{split}$ (2)
where $\delta_{a>0}$ is an indicator of a non-baseline action. By incorporating both the linear predictor $x(s,a)^{\top}\theta_{i,t}\delta_{a>0}$ and the nonlinear baseline $g_{t}(s)$, we can maintain interpretability of the action effects while also allowing flexibility in specifying the observed rewards. The individual rewards are then realized according to the model
$R_{i,t}=x(s,a)^{\top}\theta_{i,t}\delta_{a>0}+g_{t}(s)+\epsilon_{i,t},$
where we make the following standard (Abeille & Lazaric, 2017) assumption regarding $\epsilon_{i,t}$:
Assumption 2.
$\epsilon_{i,t}$ is conditionally mean zero (i.e., $\mathbb{E}[\epsilon_{i,t}|\mathcal{H}_{i,t}]=0$) and sub-Gaussian with variance $\sigma^{2}$: $\mathbb{E}[\exp(\eta\epsilon_{i,t})|\mathcal{H}_{i,t}]\leq\exp(\eta^{2}\sigma^{2}/2)$ for $\eta>0$.
Under this model assumption, the IPW estimator of Greenewald et al. (2017) produces an unbiased estimate of $\Delta_{i,t}(s,a)$; however, it suffers from high variance, which leads to suboptimal performance in practice. Our proposed algorithm DML-TS-NNR, however, relies on an improved estimator, $\tilde{R}_{i,t}^{f}(s,\bar{a})$, that is also unbiased but achieves lower variance by denoising the rewards according to a working model, $f_{i,t}(s,a)$, for the conditional mean $r_{i,t}(s,a):=\mathbb{E}\left[R_{i,t}|S_{i,t}=s,A_{i,t}=a\right]$:
[…]
where $\bar{a}$ is a pre-selected non-baseline arm; i.e., we condition on $A_{i,t}\in\{0,\bar{a}\}$. The first component is a model-based prediction of the differential reward, and the second component is a debiasing term that uses IPW to correct for potential errors in specifying $f$. This improved estimator is a form of pseudo-outcome, a common technique in the causal inference literature (Bang & Robins, 2005; Kennedy, 2020; Nie & Wager, 2021). We refer to it as a Doubly Robust Differential Reward because it produces an unbiased estimate of $\Delta_{i,t}(s,a)$ as long as either $\pi_{i,t}$ or $f_{i,t}$ is correctly specified (see proof in Appendix F.1); though, in practice, the benefit stems from (a) the robustness to misspecification of $f$ (because $\pi_{i,t}$ is known) and (b) the variance reduction introduced by de-noising the outcomes. Lemma 5 and Remark 2 in Appendix F.2 provide a proof and discussion of the variance reduction. In the following, we abbreviate $\tilde{R}_{i,t}^{f}(s,a)$ as $\tilde{R}_{i,t}^{f}$, with the state and action implied.
4 Algorithm Components
We now construct $\tilde{R}_{i,t}^{f}$, introduce the weighted least-squares loss function used by DML-TS-NNR to estimate parameters, and use nearest neighbor regularization to pool information across users.
4.1 Debiased Machine Learning
To construct $\tilde{R}_{i,t}^{f}$, we fit a working model $f_{i,t}$ for the conditional mean reward via the DML framework, which requires three main conditions: (a) Neyman Orthogonality - making the pseudo-outcome $\tilde{R}_{i,t}^{f}$ robust to misspecification of $f_{i,t}$, (b) sufficiently fast convergence of the working conditional mean model $f_{i,t}$, and (c) independence of $f_{i,t}$ and the observed outcomes. (a) is ensured by the structure of the doubly-robust differential reward. (b) requires us to introduce the following assumption:
Assumption 3.
The working model, $f_{i,t}$, satisfies $\mathbb{E}_{p(s,\bar{a})}\left[(r_{i,t}(s,\bar{a})-f_{i,t}(s,\bar{a}))^{2}\right]=o_{P}(k^{-1/2})$ and $\mathbb{E}_{p(s)}\left[(r_{i,t}(s,0)-f_{i,t}(s,0))^{2}\right]=o_{P}(k^{-1/2})$.
Examples of supervised learning methods that converge at this rate (or faster) include parametric models of fixed dimension, local mean smoothers with small dimension (Wasserman, 2006, Chapter 5), and random forests and feedforward neural networks under certain regularity conditions (Biau, 2012; Farrell et al., 2021). Because of the Neyman Orthogonality, we can achieve sub-linear regret even if $f_{i,t}$ is misspecified. Assumption 3 guarantees better asymptotic variance than action-centering (AC) bandits Greenewald et al. (2017). Without it, the worst case performance (i.e., extreme model misspecification) can lead infinite asymptotic variance and perform empirically much worse than AC bandits (Bang & Robins, 2005). In mHealth contexts, we believe the assumption is fairly weak and plausible. The final requirement of independence, (c), can be accomplished via sample-splitting techniques. Below, we present two such options.
Option 1: The first option exploits the fact that outcomes across different users are independent. Consequently, we can perform sample splitting as follows. Step 1: Randomly assign each user $i$ to one of $J$ folds. Let $I_{j}$ denote the set containing user indexes for the $j$-th fold and let $I_{j}^{\complement}$ denote its complement. Step 2: For each fold, use a supervised learning algorithm to estimate the working model for $r_{i,t}(s,a)$ denoted $\hat{f}_{i,t}^{(j)}(s,a)$ using $I_{j}^{\complement}$. Step 3: Construct the pseudo-outcomes using (3).
Option 2: Because we do not expect an adversarial environment in mHealth, we may be willing to assume that the errors, $\epsilon_{i,t}$, are independent across time. In this case, we can adapt the procedure given above by randomly assigning data to folds at the level of $(i,t)$. In this case, $I_{j}$, contains the $(i,t)$ pairs assigned to fold $j$. This option is more powerful than option 1 in that it enables us to learn heterogeneity across users in the nuisance function, $f_{i,t}$, leading to greater variance reduction.
Lastly, we note two strategies that can reduce the computational demands of sample splitting. The first is to employ certain stable estimators that do not require sample splitting, e.g., bagged estimators using subsampling, as discussed in Chen et al. (2022). The second is to fit $f_{i,t}$ in an online fashion.
4.2 Estimation via Weighted Least Squares
Having formed $\tilde{R}_{i,t}^{f}(s,\bar{a})$, we now explain how to estimate the parameters, $\theta_{i,t}\in\mathbb{R}^{p}$, determining the differential reward, $\Delta_{i,t}(s,a)$. We assume that $\theta_{i,t}$ consists of three components:
Assumption 4.
$\theta_{i,t}=\theta^{\text{shared}}+\theta^{\text{user}}_{i}+\theta^{\text{time}}_{t}$ for all $i,t$.
The parameter $\theta^{\text{shared}}$ is a “structural parameter” shared across all $i,t$. In contrast, $\theta^{\text{user}}_{i}$ and $\theta^{\text{time}}_{t}$ are “incidental parameters” specific to a given user and time point, respectively. We place these parameters in a single column vector, $\theta$, as $\textrm{vec}(\theta^{\text{shared}},\theta^{\text{user}}_{1},\ldots,\theta^{\text{user}}_{K},\theta^{\text{time}}_{1},\ldots,\theta^{\text{time}}_{K})$, where $K$ is the total number of stages. We then form a corresponding feature vector, $\phi_{i,t}=\phi(x_{i,t})$, where $\phi$ inserts $x_{i,t}$ into the positions corresponding to $\theta^{\text{shared}}$, $\theta^{\text{user}}_{i}$, and $\theta^{\text{time}}_{t}$ with zeros in all other locations; i.e., letting $p:=\text{dim}(x_{i,t})$, and the feature vector $\phi(x_{i,t})\in\mathbb{R}^{(2K+1)p}$ is defined as follows:
[…]
We could then estimate $\theta_{i,t}$ via a weighted least squares (WLS) regression that minimizes the following objective function:
$\ell_{\text{WLS}}(\theta)=\sum_{i=1}^{n}\sum_{t=1}^{K-i+1}\tilde{\sigma}_{i,t}^{2}\big{(}\tilde{R}_{i,t}^{f}-\phi_{i,t}^{\top}\theta\big{)}^{2},$ (5)
where $\tilde{\sigma}_{i,t}^{2}=\pi_{i,t}(0|S_{i,t})\cdot(1-\pi_{i,t}(0|S_{i,t}))$. However, $\ell_{\text{WLS}}(\theta)$ is over-parameterized and does not leverage the available network information. The next section modifies $\ell_{\text{WLS}}(\theta)$ by adding regularization, producing the final loss function used in DML-TS-NNR.
4.3 Nearest-Neighbor Regularization
We assume access to network information relating neighboring users and time points. These networks are defined by graphs $G_{\text{user}}=(V_{\text{user}},E_{\text{user}})$ and $G_{\text{time}}=(V_{\text{time}},E_{\text{time}})$, where $V$ denotes vertices and $E$, edges. In practice, these networks can be formed via known clusters (e.g., shared schools, companies, geographic locations) or similar covariates. In the case of $G_{\text{time}}$, a reasonable choice is to construct a graph of neighboring sequential time points; i.e., $t=1\leftrightarrow t=2\leftrightarrow t=3\dots$. We assume that neighboring units have similar incidental parameters as follows:
Assumption 5.
There exists $D_{\text{user}},D_{\text{time}}\in\mathbb{R}^{+}$ such that $\|\theta^{\text{user}}_{i}-\theta^{\text{user}}_{j}\|_{2}^{2}\leq D_{\text{user}}$ and $\|\theta^{\text{time}}_{i}-\theta^{\text{time}}_{j}\|_{2}^{2}\leq D_{\text{time}}$, for all $i,j$.
Assumption 5 suggests we may be able to improve our estimates of $\theta$ by regularizing neighboring values of $\theta^{\text{user}}_{i}$ and $\theta^{\text{time}}_{i}$ toward each other. This can be accomplished by using an $L^{2}$ Laplacian penalty. We illustrate the idea using $G_{\text{user}}$; the application to $G_{\text{time}}$ is similar.
First, we form the incidence matrix, $Q\in\mathbb{R}^{N\times N}$, with the entry $Q_{v,e}$ corresponding to the $v$-th vertex (user) and $e$-th edge. Denote the vertices of $e$ as $v_{i}$ and $v_{j}$ with $i>j$. $Q_{v,e}$ is then equal to 1 if $v=v_{i}$, -1 if $v=v_{j}$, and 0 otherwise. The Laplacian matrix is then defined as $L=QQ^{\top}$. Similar to Yang et al. (2020), we can then form a network cohesion penalty across users as follows:
$\text{tr}(\Theta_{{\text{user}}}^{\top}L_{\text{user}}\Theta_{{\text{user}}})=\sum_{(i,j)\in E_{\text{user}}}\|\theta^{{\text{user}}}_{i}-\theta^{{\text{user}}}_{j}\|_{2}^{2},$
where $\Theta_{\text{user}}:=(\theta^{\text{user}}_{1},\ldots,\theta^{\text{user}}_{N})^{\top}\in\mathbb{R}^{N\times p}$ and $L_{\text{user}}$ is the Laplacian matrix for users. The penalty is small when $\theta^{\text{user}}_{i}$ and $\theta^{\text{user}}_{j}$ are close for connected users. We employ a shared regularization hyperparameter $\lambda$, for $G_{\text{user}}$ and $G_{\text{time}}$, and include a standard $L^{2}$ penalty on $\theta$ with hyperparameter $\gamma$. The full penalization matrix $V_{0}\in\mathbb{R}^{(2K+1)p\times(2K+1)p}$ is
$V_{0}=\operatorname{diag}(\gamma I_{p},\lambda L^{\text{user}}_{\otimes}+\gamma I_{Kp},\lambda L^{\text{time}}_{\otimes}+\gamma I_{Kp}),$ (6)
where $L^{\text{user}}_{\otimes}:=L_{\text{user}}\otimes I_{p}$ , the Kronecker product of $L_{\text{user}}$ with $I_{p}$. We then adapt $\ell_{\text{WLS}}(\theta)$ to include a corresponding penalty term:
$\ell(\theta)=\sum_{i=1}^{n}\sum_{t=1}^{K-i+1}\tilde{\sigma}_{i,t}^{2}\big{(}\tilde{R}_{i,t}^{f}-\phi_{i,t}^{\top}\theta\big{)}^{2}+\theta^{\top}V_{0}\theta.$ (7)
In Section 5 we show that (7) leads to a Thompson sampling algorithm for action selection.
5 Thompson Sampling Algorithm
We now describe the Thompson sampling procedure and provide a high-probability regret bound.
5.1 Algorithm Description
The minimizer of the weighted regularized least squares loss in (7) can be expressed as $\hat{\theta}=V^{-1}b$, where
$V=V_{0}+\sum_{i=1}^{n}\sum_{t=1}^{K-i+1}\tilde{\sigma}_{i,t}^{2}\phi_{i,t}\phi_{i,t}^{\top},\quad b=\sum_{i=1}^{n}\sum_{t=1}^{K-i+1}\tilde{R}_{i,t}^{f}\sigma_{i,t}^{2}\phi_{i,t}.$
In analogy to Bayesian linear regression, the penalized and weighted Gram matrix, $V$, plays the role of a posterior precision matrix. This motivates an algorithm in which we randomly perturb the point estimate, $\hat{\theta}$, scaling by an inverse square-root matrix, $V^{-1/2}$, times a mean-zero random deviate, $\eta$. The perturbed estimate, $\tilde{\theta}$, can be written as $\tilde{\theta}=\hat{\theta}+\beta(\delta)V^{-1/2}\eta,$ where $\beta(\delta)$ is defined as
$\beta(\delta)$ $=v\left[2\log\left(\frac{\det(V)^{1/2}}{\det(V)^{1/2}\delta/2}\right)\right]^{1/2}+B,$ (8)
$B$ $=K\frac{\lambda}{\sqrt{\gamma}}\left(D_{\text{user}}+D_{\text{time}}\right)+\sqrt{\gamma n}\left(B_{\text{user}}+B_{\text{time}}\right).$ (9)
The value $v$ is defined in Corollary 4 and can be set to a sufficiently large constant. The hyperparameter $\delta$ determines the probability with which the regret bound holds. The random deviate, $\eta$, is drawn from a distribution, $\mathcal{D}^{TS}$, satisfying the technical conditions given in Definition 1 in Appendix E. In practice, we recommend setting $\mathcal{D}^{TS}$ to a multivariate Gaussian distribution or t-distribution to simplify calculations. Given $\tilde{\theta}$, we then compute $\bar{A}_{i,t}=\operatorname{argmax}_{a\in[q]\backslash\{0\}}\phi_{i,t}^{\top}\tilde{\theta}$ and randomly select $A_{i,t}=0$ with probability
$\pi_{i,t}(0|\bar{A}_{i,t}):=\max\big{(}\pi_{\min},\min\{\pi_{\max},\Pr(\phi_{i,t}^{\top}\tilde{\theta}>0)\}\big{)},$ (10)
or $A_{i,t}=\bar{A}_{i,t}$ with probability $1-\pi_{i,t}(0|\bar{A}_{i,t})$. The action selection is summarized in Algorithm 1.
Algorithm 1 Action selection
0:  $V$, $b$, $\pi_{\min}$, $\pi_{\max}$, $\phi_{i,t}$, $\beta(\delta)$
  $\hat{\theta}=V^{-1}b$
  Sample $\eta\sim\mathcal{D}^{TS}$
  $\tilde{\theta}=\hat{\theta}+\beta(\delta)V^{-1/2}\eta$, with $\beta(\delta)$ given in (8)
  $\bar{A}_{i,t}=\operatorname{argmax}_{a\in[q]\backslash\{0\}}\phi_{i,t}^{\top}\tilde{\theta}$
  Calculate $\pi_{i,t}(0|\bar{A}_{i,t})$ according to (10)
  With probability $\pi_{i,t}(0|\bar{A}_{i,t})$, play action 00; otherwise, play action $\bar{A}_{i,t}$
Having detailed the action selection algorithm, we now summarize DML-TS-NNR in Algorithm 2. The algorithm includes an outer loop that iterates through the stages from $k=1$ to $k=K$. Within each stage, we handle users sequentially, iteratively selecting actions according to Algorithm 1, observing rewards, and updating $V$ and $b$ accordingly.
Algorithm 2 DML-TS-NNR
  Initialize $V_{0}$ according to (6)
  $V\leftarrow V_{0}$
  $b\leftarrow 0$
  for $k=1,\ldots,K$  do
     for $(i,j)=(1,k),(2,k-1),\dots,(k,1)$  do
        Choose an arm according to Algorithm 1 using $V$ and $b$
        Observe rewards, $R_{i,t}$
        Construct pseudo-rewards, $\tilde{R}_{i,t}^{f^{(m)}}$, as explained in Section 4.1
        Set $\phi=\phi(x_{i,t})$ as described in (4)
        Update weighted Gram matrix: $V\leftarrow V+\tilde{\sigma}^{2}_{i,t}\phi_{i,t}\phi_{i,t}^{\top}$
        Update feature-outcome products: $b\leftarrow b+\tilde{\sigma}_{i,t}^{2}\tilde{R}_{i,t}^{f}\phi_{i,t}$
     end for
  end for
5.2 Regret bound
We first define regret, $\boldsymbol{\operatorname{Regret}}_{K}$, in terms of the following average across stages:
$\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\big{[}\pi_{i,t}^{\star}(\bar{A}_{i,t}^{*}|S_{i,t})\cdot x(S_{i,t},\bar{A}_{i,t}^{*})^{\top}\theta^{*}_{i,t}\\ -\pi_{i,t}(\bar{A}_{i,t}|S_{i,t})\cdot x(S_{i,t},\bar{A}_{i,t})^{\top}\theta^{*}_{i,t}\big{]}.$
The inner sum calculates regret (the difference between the expected reward of the optimal action and the chosen action) weighted by the probability of the action, following Greenewald et al. (2017), for all individuals at stage $k$. We then average these values to account for the growing number of pairs $(i,t)$ as the stages progress. The cumulative regret is the sum of these averaged values over all stages.
The rewards associated with the baseline action are excluded since $\Delta_{i,t}(S_{i,t},0)=0$. $\boldsymbol{\operatorname{Regret}}_{K}$ counts as a pseudo-regret because the randomness attributed to $\{\epsilon_{i,t}\}_{t=1}^{K}$ has been eliminated (Audibert et al., 2003). Before presenting the regret bound, we introduce one more assumption:
Assumption 6.
$\|x(s,a)\|\leq 1$ for all $s$ and $a$. Also, there exists a known value $M\in\mathbb{R}^{+}$ bounding both $\|\theta_{i,t}\|$ and $|g_{t}(s)|$ for all $i$, $t$, and $s$. Further, $|f_{i,t}|\leq 2M$.
Theorem 1.
Under Assumptions 16, with probability at least $1-\delta$, $\boldsymbol{\operatorname{Regret}}_{K}$ satisfies
$O\left(\left(\beta_{K}(\delta^{\prime})+\gamma_{K}(\delta^{\prime})\left[1+\frac{4}{p}\right]\right)\boldsymbol{\cdot}\log(K)\sqrt{4cKp}\right)$
where $\delta^{\prime}=\delta/4K$ and $\min(\pi(0|s),1-\pi(0|s))>1/c$. The exact regret bound is given and both $\beta_{K}$ and $\gamma_{K}$ are defined in Appendix F.
Theorem 1’s proof is in Appendix F. It resembles  Abeille & Lazaric (2017), albeit with adjustments to address (a) the nonlinear baseline and (b) the increasing pool of users and time points. The non-linear baseline requires deriving the new sub-Gaussian variance factor for the centered pseudo-reward and adapting classic contextual bandit theory (Abbasi-Yadkori et al., 2011) to our weighted least squares setting. The increasing pool of users and time points requires modifying the Abeille & Lazaric (2017) proof with a careful use of stages and Cauchy Schwarz. The overall bound scales as
$O\left(p^{3/2}\sqrt{\log^{2}(K)K}\right).$
DML is particularly advantageous for $\beta_{K}(\delta^{\prime})$ and $\gamma_{K}(\delta^{\prime})$, which rely on the convergence rate of the model $f$ towards the true mean differential reward $r$, as discussed further in Appendix F.2. Another interesting aspect of the regret bound is its dependence solely on the dimension, $p$, of the differential reward model, rather than the dimension of the baseline reward. This characteristic allows us to achieve robust regret bounds even in cases where the baseline reward is highly complex.
6 Experiments
This section presents results from applying DML-TS-NNR in a simulated mHealth study with staggered recruitment and an off-policy comparison study for cardiac rehabilitation mHealth interventions. The simulations were implemented using Python and the results were generated using individual compute nodes with two 3.0 GHz Intel Xeon Gold 6154 processors and 180 GB of RAM. Case studies were implemented using R 4.2.2 and results were generated on a cluster composed of individual compute nodes with 2.10 GHz Intel Xeon Gold 6230 processors and 192 GB of RAM.
6.1 Competitor Comparison Simulation
In this section, we compare DML-TS-NNR to four competing methods in simulation. We implemented DML-TS-NNR using Option 2 with a bagged ensemble of stochastic gradient trees (Gouk et al., 2019; Mastelini et al., 2021) trained online via the River library (Montiel et al., 2021). We programmed the linear algebra operations using SuiteSparse (Davis & Hu, 2011) to take advantage of the sparse nature of the Laplacian matrices and $\phi_{i,t}$ from Algorithm 2.
Our competing baselines are (a) Standard: Standard Thompson sampling for linear contextual bandits, (b) AC: the Action-Centered contextual bandit algorithm (Greenewald et al., 2017), (c) IntelPooling: The intelligentpooling method of Tomkins et al. (2021) fixing the variance parameters close to their true values, and (d) Neural-Linear: a method that uses a pre-trained neural network to transform the feature space for the baseline reward (similar to the Neural Linear method of Riquelme et al. (2018)).
We compare these methods under three settings: Homogeneous Users, Heterogeneous Users, and Nonlinear. The first two involve a linear baseline model and time-homogeneous parameters, with the second setting having distinct user parameters. The third setting, designed to mirror the challenges of mHealth studies, includes a nonlinear baseline and both user- and time-specific parameters. For each setting, we simulate 200 stages following the staged recruitment regime depicted in Figure 3 in Appendix A, and we repeat the full 200-stage simulation 50 times. Appendix B.1 provides details on the setup and a link to our implementation.
Figure 1: Cumulative regret in the (a) Homogeneous Users, (b) Heterogeneous Users, and (c) Nonlinear settings. DML-TS-NNR performs competitively in the first setting (the simplest), and it substantially outperforms the next-best method (IntelPooling) in the others. Pairwise comparisons in Table 2 in Appendix B.3) reveal that DML-TS-NNR outperformed all competitors shown above in 49 of 50 repetitions in the Nonlinear setting (panel (c)).
Figure 1 displays the average cumulative regret for each method. DML-TS-NNR and IntelPooling have similar performance in the Homogeneous Users setting, but DML-TS-NNR excels in the Heterogeneous Users and Nonlinear settings, with significantly lower cumulative regret by stage 200. This is expected as DML-TS-NNR can utilize network information and model nonlinear baselines in these settings. However, in the Homogeneous Users setting, where no network information is available and a linear baseline is used, DML-TS-NNR does not outperform IntelPooling.
Appendices B.2, B.3, and B.4 offer more results, including other method comparisons, a rectangular data array simulation, and pairwise statistical comparisons. The latter show DML-TS-NNR outperformed all competitors in at least 49 of 50 repetitions in the Nonlinear setting, indicating even higher statistical confidence than Figure 1 suggests.
6.2 Valentine Study Analysis Results
Figure 2: (left) Boxplot of unbiased estimates of the average per-trial reward for all five competing algorithms, relative to the reward obtained under the pre-specified Valentine randomization policy across 100 bootstrap samples. Within each box, the asterisk ($\ast$) indicates the mean value, while the mid-bar represents the median. (right) Heatmap of p-values from the pairwise paired t-tests.
In this section, we compare DML-TS-NNR to the above algorithms via off-policy evaluation using data from the Valentine Study (Jeganathan et al., 2022), a prospective, randomized-controlled, remotely-administered trial designed to evaluate an mHealth intervention to supplement cardiac rehabilitation for low- and moderate-risk patients. In the analyzed dataset, participants were randomized to receive or not receive contextually tailored notifications promoting low-level physical activity and exercise throughout the day. The left of Figure 2 shows the estimated improvement in the average reward over the original constant randomization, averaged over stages (K = 120) and participants (N=108). We see that DML-TS-NNR achieved the highest average reward; its average performance (across bootstrap replications) is higher than the 75th percentile of all other methods.
To further analyze the improvement of DML-TS-NNR relative to the other methods, we aim to test whether the proposed algorithm consistently and significantly improves cumulative rewards across all assessed datasets. Here we conducted pairwise paired t-tests with one-sided alternative hypotheses. The null hypothesis ($H_{0}$) for these tests is that the two algorithms being compared achieve the same average reward. The alternative hypothesis ($H_{1}$) is that the algorithm listed in the column achieves higher average rewards than the algorithm listed in the row. The right of Figure 2 displays the p-values obtained from these pairwise t-tests. The last column’s dark shade indicates that the proposed DML-TS-NNR algorithm achieves significantly higher rewards than the other five competing algorithms ($p\leq 0.01$ for all pairwise comparisons).. This implies that after implementing our proposed algorithm, the step counts increased by $3.5\%$ more than what the constant randomization policy achieved. This translates to an increase of 140 steps per day in step count, given a baseline of 1000 steps per hour. See Appendix C for additional implementation details.
We performed an additional off-policy comparison using data from the Intern Health Study (IHS) (NeCamp et al., 2020), as shown in Figure 10. The results further demonstrate the competitive performance of our DML-TS-NNR algorithm, with the AC algorithm also showing comparable performance in this dataset. Further details on the analysis can be found in Appendix D.
7 Discussion
This paper introduces DML-TS-NNR, a new contextual bandit algorithm for mHealth studies, using DML and network cohesion penalties to model complex rewards and pool information across individuals and time. This enhances statistical precision, enabling rapid learning of accurate, context-specific mHealth treatment policies and improved rewards.
While DML-TS-NNR surpasses current methods, it has four potential improvement areas: it only considers immediate rewards, ignoring treatment fatigue; its computation of log-determinants and matrix inverses is expensive for large matrices; it assumes linear differential rewards and a known network with binary edges; and it requires specification of several hyperparameters. Future work will address these issues.
Appendix A Recruitment Regime Illustration
Figure 3: Illustration of the staged recruitment scheme. At each recruitment stage (each time point), a new participant is recruited and observed; at the same time, all participants who were recruited prior to the current stage are also observed again. Observations are not collected from participants who have yet to be recruited. For simplicity, we assume one participant is recruited at each stage.
Appendix B Additional Details for Simulation Study
This appendix provides details regarding the setup of our simulation study, additional results not presented in the main paper, and results from additional simulations employing different competitors and/or simulation designs. We summarize the differences between the competing algorithms in Table 1.
User-specific Parameters Time-specific parameters Pools Across Users Enforces User Network Cohesion Enforces Time Network Cohesion Allows Nonlinear Baseline Nonlinear Baseline Model
DML-TS-NNR $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
DML-TS-NNR-BLM $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\times$
DML-TS-SU $\times$ $\surd$ $\surd$ N/A $\surd$ $\surd$ $\surd$
NNR-Linear $\surd$ $\times$ $\surd$ $\surd$ N/A $\times$ $\times$
IntelPooling $\surd$ $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\times$
Neural-Linear $\surd$ $\times$ $\times$ N/A N/A $\surd$ $\surd$
Standard $\surd$ $\times$ $\times$ N/A N/A $\times$ $\times$
AC $\surd$ $\times$ $\times$ N/A N/A $\surd$ $\times$
Table 1: This table compares the methods tested in the simulation study based on the design components listed in the columns, with $\surd$ indicating that the method includes the component listed in the column, $\times$ indicating that it does not, and N/A indicating that the component is not applicable. This table can be used to investigate the effect of certain model components on cumulative regret; for example, comparing DML-TS-NNR and DML-TS-NNR-BLM helps us understand the impact of modeling the nonlinear baseline—the only column that differs between these two methods.
B.1 Setup Details
The code for the simulation study is fully containerized and publicly available at https://github.com/anonymous-conf-submitter-896697/mixed-effects-bandit. The simulation study involves a generative model of the following form for user $i$ at time $t$:
$R_{it}=g(S_{it})+x(S_{it},A_{it})^{{\top}}\,\theta_{it}+\epsilon_{it},\quad\epsilon_{it}\sim\mathcal{N}(0,1)$
Here $S_{it}=(s_{1},s_{2})\in\mathbb{R}^{2}$ is a context vector, with both dimensions $\overset{iid}{\sim}U(-1,1)$. We set $x(s,a)=a\,(1,s_{1},s_{2})$. For simplicity, we set $g$ to a time-homogeneous function. The specific nature of the function varies across the following three settings mentioned in Section 6.1:
  • Homogeneous Users: Standard contextual bandit assumptions with a linear baseline and no user- or time-specific parameters. The linear baseline is $g(S_{it})=2-2s_{1}+3s_{2}$, and the causal parameter is $\theta_{it}=(1,0.5,-4)$ such that the optimal action varies across the state space.
  • Heterogeneous Users: Same as the above but each user’s causal parameter has iid $\mathcal{N}(0,1)$ noise added to it.
  • Nonlinear: The general setting discussed in the paper with a nonlinear baseline, user-specific parameters, and time-specific parameters. The base causal parameter and user-specific parameters are the same as in the previous two settings. The nonlinear baseline and time-specific parameter are shown in Figure 4.
Figure 4: (left) The baseline reward function $g(S_{it})$ used in the simulation study. The proposed method allows this function to be a nonlinear function of the context vectors. The baseline was generated using a combination of recursive partitioning and by summing scaled, shifted, and rotated Gaussian densities. (right) The time-specific parameters used in the simulation study. These parameters cause the advantage function to vary over time. We set them such that the advantage function changes quickly at the beginning of the study then stabilizes.
We assume that the data are observed via a staged recruitment scheme, as illustrated in Figure 3 in Appendix A. For computational convenience, we update parameters and select actions in batches. If, for instance, we observe twenty users at a given stage, we update our estimates of the relevant causal parameters and select actions for all twenty users simultaneously. This strategy offers a slight computational advantage with limited implications in terms of statistical performance.
For simplicity, we assume that the nearest-neighbor network is known and set the relevant hyperparameters accordingly. We took care to set the hyperparameters such that DML-TS-NNR performs a similar amount of shrinkage compared to other methods, especially IntelPooling, which effectively uses a separate penalty matrix for users and time. To accomplish this, we used the same penalty matrices for DML-TS-NNR and IntelPooling, effectively generalizing the scalar $\gamma$ penalty discussed in the main paper. For the purposes of calculating the constant $B$ given in Algorithm 2, we used a separate value of $\gamma$ for both users and time and set them equal to the maximum eigenvalue of the corresponding penalty matrix. We use 5 neighbors within the DML methods and set the other hyperparameters as follows: $\lambda=1$, and $\delta=0.01$, and $v_{k}=1$ (i.e., $v_{k}$ is constant).
For the Neural-Linear method, we generate a $200\times 200$ array of baseline rewards to train the neural network prior to running the bandit algorithm. Consequently, the results shown in the paper for Neural-Linear are better than would be observed in practice because we allowed the Neural-Linear method to leverage data that we did not make available to the other methods. This setup offers the computational benefit of not needing to update the neural network within bandit replications, which substantially reduces the necessary computation time.
Aside from the input features used, the Neural-Linear method has the same implementation as the Standard method. The Neural-Linear method uses the output from the last hidden layer of a neural network to model the baseline reward. However, we use the original features (the state vectors) to model the advantage function because the true advantage function is, in fact, linear in these features.
Our neural networks consisted of four hidden layers with 10, 20, 20, and 10 nodes, respectively. The first two employ the ReLU activation function (Nair & Hinton, 2010) while the latter two employ the hyperbolic tangent. We chose to use the hyperbolic tangent for the last two layers because Snoek et al. (2015) found that smooth activation functions such as the hyperbolic tangent were advantageous in their neural bandit algorithm. The loss function was the mean squared error between the neural networks’ output and the baseline reward on a simulated data set. We trained our networks using the Adam optimizer (Kingma & Ba, 2014) with batch sizes of 200 for between 20 and 50 epochs. We simulated a separate validation data set to check that our model had converged and was generating accurate predictions.
Figure 5 compares the true baseline reward function in the nonlinear setting (left) to that estimated by the neural network (right). We see that the neural network produced an accurate approximation of the baseline reward, which helps explain the good performance of the Neural-Linear method relative to other baseline approaches, such as Standard.
Figure 5: (left) The baseline reward function $g(S_{it})$ used in the simulation study compared to (right) the estimated baseline reward from our neural network in the nonlinear setting.
We include the Neural-Linear method primarily to demonstrate that correctly modeling the baseline is not sufficient to ensure good performance. In the mobile health context, algorithms should also be able to (1) efficiently pool data across users and time and (2) leverage network information. The Neural-Linear method satisfies neither of these criteria. Note that a neural network could be used to model the baseline rewards as part of our algorithm. Future work could consider allowing the differential rewards themselves to also be complex nonlinear functions, which could be accomplished by combining our method with Neural-Linear. We leave the details to future work.
B.2 Additional Method Comparisons
Our results in the main paper indicate that DML-TS-NNR outperforms existing methods, especially in the complex longitudinal setting that we are targeting. These positive results beg the question: What aspects of DML-TS-NNR contribute most to its superior performance?
In this section, we describe an additional simulation study designed to answer this question. In it, we compare DML-TS-NNR (as implemented in the main paper) to three additional comparison methods: DML-TS-NNR-BLM, DML-TS-SU, and NNR-Linear. Each comparison isolates particular aspects of our algorithm to understand its contribution to DML-TS-NNR’s performance.
DML-TS-NNR-BLM is a version of DML-TS-NNR that employs Bagged Linear Models as the baseline supervised learning algorithm. Unlike the method implemented in the main paper (which uses bagged stochastic gradient trees), this baseline model is misspecified in the nonlinear setting; i.e., it cannot accurately model the nonlinear baseline reward function. Consequently, this comparison will allow us to gauge the impact of the supervised learning method used in our method. In particular, it should allow us to validate whether the algorithm is robust to baseline misspecification as the theory suggests.
DML-TS-SU is a Single-User version of our algorithm. Rather than having a separate $\theta_{i}$ for each user, this method estimates a single advantage function that is shared across all users. It does, however, employ DML, time effects, and nearest-neighbor regularization (for the time effects). Consequently, this comparison will help us understand the impact of the user-specific parameters in our simulation.
NNR-Linear is similar to the existing network-cohesion bandit algorithms described in the main paper. It includes a distinct $\theta_{i}$ for each user and regularizes them toward each other using a Laplacian penalty. It does not, however, use DML or time effects. As a result, this comparison isolates the impact of these two factors and should provide evidence that DML-TS-NNR is more broadly applicable than existing network-cohesion bandit algorithms.
Figure 6 displays the cumulative regret as a function of the stage for these methods across the three settings used in the main paper. We see that all methods perform similarly in panel (a), the setting with homogeneous users. However, in panel (b) we see that the regret achieved by DML-TS-SU begins to resemble a straight upward-sloping line because it cannot appropriately model the user-specific advantage functions.
In panel (c), we observe a much wider spread of performance, with DML-TS-NNR and DML-TS-NNR-BLM performing similarly and substantially outperforming the other methods. The fact that DML-TS-NNR and DML-TS-NNR-BLM perform similarly indicates that our algorithm is, in fact, robust to misspecification as expected. The comparison to NNR-Linear reveals that the DML and time effects—the new components of our method compared to existing network-cohesion bandit algorithms—do contribute meaningfully to DML-TS-NNR’s superior performance. We found in additional simulations studies not reported here that the gap between NNR-Linear and our method widens as the magnitude of the time effects increases.
Figure 6: Cumulative regret in the (a) Homogeneous Users, (b) Heterogeneous Users, and (c) Nonlinear settings for the additional comparison methods. The fact that DML-TS-NNR outperforms DML-TS-SU demonstrates the importance of the user-specific parameters. The fact that DML-TS-NNR outperforms NNR-Linear shows that previous network cohesion approaches do not adequately address the (1) nonlinear nature of the baseline reward and (2) presence of time effects. The comparable performance of DML-TS-NNR-BLM to DML-TS-NNR highlights how our algorithm is robust to misspecification of the baseline reward model.
B.3 Pairwise Comparisons
Table 2 shows pairwise comparisons between methods across the three settings. The individual cells indicate the percentage of repetitions (out of 50) in which the method listed in the row outperformed the method listed in the column. The asterisks indicate p-values below 0.05 from paired two-sided t-tests on the differences in final regret. The Avg column indicates the average pairwise win percentage.
DML-TS-NNR and DML-TS-NNR-BLM perform well across all three settings and the difference between them is statistically indistinguishable in 2 of 3 settings. These methods perform about equally well compared to IntelPooling and NNR-Linear in the first setting, but they perform better in the other two. In the Nonlinear setting in particular, DML-TS-NNR and DML-TS-NNR-BLM substantially outperform all other methods, and the pairwise differences are statistically significant at the 5% level.
Homogeneous Users
1 2 3 4 5 6 7 8 Avg
1. DML-TS-NNR - 58%* 48% 52% 32%* 100%* 100%* 100%* 70%
2. DML-TS-NNR-BLM 42%* - 40%* 36%* 22%* 100%* 100%* 100%* 63%
3. DML-TS-SU 52% 60%* - 48% 34%* 100%* 100%* 100%* 71%
4. NNR-Linear 48% 64%* 52% - 32%* 100%* 100%* 100%* 71%
5. IntelPooling 68%* 78%* 66%* 68%* - 100%* 100%* 100%* 83%
6. Neural-Linear 0%* 0%* 0%* 0%* 0%* - 100%* 100%* 29%
7. Standard 0%* 0%* 0%* 0%* 0%* 0%* - 100%* 14%
8. AC 0%* 0%* 0%* 0%* 0%* 0%* 0%* - 0%
Heterogeneous Users
1 2 3 4 5 6 7 8 Avg
1. DML-TS-NNR - 46% 64% 58% 90%* 100%* 100%* 100%* 80%
2. DML-TS-NNR-BLM 54% - 64%* 60% 100%* 100%* 100%* 100%* 83%
3. DML-TS-SU 36% 36%* - 44% 88%* 100%* 100%* 100%* 72%
4. NNR-Linear 42% 40% 56% - 94%* 100%* 100%* 100%* 76%
5. IntelPooling 10%* 0%* 12%* 6%* - 80%* 100%* 100%* 44%
6. Neural-Linear 0%* 0%* 0%* 0%* 20%* - 100%* 100%* 31%
7. Standard 0%* 0%* 0%* 0%* 0%* 0%* - 100%* 14%
8. AC 0%* 0%* 0%* 0%* 0%* 0%* 0%* - 0%
Nonlinear
1 2 3 4 5 6 7 8 Avg
1. DML-TS-NNR - 54% 100%* 98%* 98%* 100%* 100%* 100%* 93%
2. DML-TS-NNR-BLM 46% - 94%* 98%* 100%* 100%* 100%* 100%* 91%
3. DML-TS-SU 0%* 6%* - 42% 78%* 100%* 100%* 100%* 61%
4. NNR-Linear 2%* 2%* 58% - 82%* 100%* 100%* 100%* 63%
5. IntelPooling 2%* 0%* 22%* 18%* - 98%* 100%* 100%* 49%
6. Neural-Linear 0%* 0%* 0%* 0%* 2%* - 98%* 100%* 29%
7. Standard 0%* 0%* 0%* 0%* 0%* 2%* - 100%* 15%
8. AC 0%* 0%* 0%* 0%* 0%* 0%* 0%* - 0%
Table 2: Pairwise comparisons between methods in the three settings of the main simulation. Each cell indicates the percent of repetitions (out of 50) in which the method listed in the row outperformed the method listed in the column in term of final regret. Asterisks indicate p-values below 0.05 from paired two-sided t-tests on the differences in final regret. DML-TS-NNR and DML-TS-NNR-BLM perform well in all three settings, and their final regret is statistically indistinguishable in 2 of 3 settings. They substantially outperform all other methods in the Nonlinear setting.
B.4 Simulation with Rectangular Data Array
The main simulation involves simulating data from a triangular data array. At the 200-th (final) stage, the algorithm has observed 200 rewards for user 1, 199 rewards for user 2, and so on.
In this section, we simulate actions and rewards under a rectangular array with 100 users and 100 time points. Although we still follow the staged recruitment regime depicted in Figure 3 in Appendix A, at stage 100 we stop sampling actions and rewards for user 1; at stage 101 we stop sampling for user 2; and so on until we have sampled 100 time points for all 100 users. Aside from the shape of the data array, the setup is the same for this simulation as for the main simulation.
Figure 7: Cumulative regret in the (a) Homogeneous Users, (b) Heterogeneous Users, and (c) Nonlinear settings using a rectangular array of data in which we observe 100 time points for 100 users in a stagewise fashion as depicted in Figure 3 in Appendix A. Similar to Figure 1, DML-TS-NNR is competitive in the first setting and substantially outperforms the competitors in the other settings.
Figure 8: Cumulative regret in the (a) Homogeneous Users, (b) Heterogeneous Users, and (c) Nonlinear settings using a rectangular array of data in which we observe 100 time points for 100 users in a stagewise fashion as depicted in Figure 3 in Appendix A. We observe the same performance ordering as in Figure 6, but here the relative differences are even larger.
The cumulative regret for these methods as a function of time—not stage—is shown in Figure 7. Qualitatively, the results are nearly identical to those from the main simulation (compare to Figure 1). The differences in final regret are even larger in this simulation than they were in the main simulation, presumably because this study exhibits greater variation in time effects due to the balance across time.
Figure 8 displays comparisons similar to those shown in Figure 6 but for the rectangular simulation. We again see that the results are qualitatively similar to the main simulation. The methods perform similarly in the Homogeneous Users setting. DML-TS-SU performs poorly in the Heterogeneous Users setting, but the other methods perform similarly. Finally, DML-TS-NNR and DML-TS-NNR-BLM substantially outperform the other methods in the nonlinear setting. Again, the differences in performance are even larger in this simulation than in the main simulation.
Homogeneous Users
1 2 3 4 5 6 7 8 Avg
1. DML-TS-NNR - 54% 46% 42% 40% 100%* 100%* 100%* 69%
2. DML-TS-NNR-BLM 46% - 46% 38%* 32%* 100%* 100%* 100%* 66%
3. DML-TS-SU 54% 54% - 44% 34% 100%* 100%* 100%* 69%
4. NNR-Linear 58% 62%* 56% - 52% 100%* 100%* 100%* 75%
5. IntelPooling 60% 68%* 66% 48% - 100%* 100%* 100%* 77%
6. Neural-Linear 0%* 0%* 0%* 0%* 0%* - 100%* 100%* 29%
7. Standard 0%* 0%* 0%* 0%* 0%* 0%* - 100%* 14%
8. AC 0%* 0%* 0%* 0%* 0%* 0%* 0%* - 0%
Heterogeneous Users
1 2 3 4 5 6 7 8 Avg
1. DML-TS-NNR - 56% 100%* 44% 100%* 100%* 100%* 100%* 86%
2. DML-TS-NNR-BLM 44% - 100%* 38% 100%* 100%* 100%* 100%* 83%
3. DML-TS-SU 0%* 0%* - 2%* 92%* 100%* 100%* 100%* 56%
4. NNR-Linear 56% 62% 98%* - 100%* 100%* 100%* 100%* 88%
5. IntelPooling 0%* 0%* 8%* 0%* - 96%* 100%* 100%* 43%
6. Neural-Linear 0%* 0%* 0%* 0%* 4%* - 100%* 100%* 29%
7. Standard 0%* 0%* 0%* 0%* 0%* 0%* - 100%* 14%
8. AC 0%* 0%* 0%* 0%* 0%* 0%* 0%* - 0%
Nonlinear
1 2 3 4 5 6 7 8 Avg
1. DML-TS-NNR - 60%* 100%* 100%* 100%* 100%* 100%* 100%* 94%
2. DML-TS-NNR-BLM 40%* - 100%* 100%* 100%* 100%* 100%* 100%* 91%
3. DML-TS-SU 0%* 0%* - 28%* 78%* 100%* 100%* 100%* 58%
4. NNR-Linear 0%* 0%* 72%* - 86%* 100%* 100%* 100%* 65%
5. IntelPooling 0%* 0%* 22%* 14%* - 98%* 100%* 100%* 48%
6. Neural-Linear 0%* 0%* 0%* 0%* 2%* - 100%* 100%* 29%
7. Standard 0%* 0%* 0%* 0%* 0%* 0%* - 100%* 14%
8. AC 0%* 0%* 0%* 0%* 0%* 0%* 0%* - 0%
Table 3: Pairwise comparisons between methods in the three settings of the simulation with a rectangular array of data. As in Table 2, each cell indicates the percent of repetitions (out of 50) in which the method listed in the row outperformed the method listed in the column in term of final regret. Asterisks indicate p-values below 0.05 from paired two-sided t-tests on the differences in final regret. The qualitative results are similar to those of Table 2. DML-TS-NNR and DML-TS-NNR-BLM perform well across all three settings and substantially outperform all other methods in the Nonlinear setting.
Table 3 displays results from pairwise method comparisons, similar to those shown in Table 2. Again the results are qualitatively similar to but slightly more exaggerated than those from the main simulation. In particular, in this simulation study, DML-TS-NNR and DML-TS-NNR-BLM outperform all other methods in 100% of repetitions in the nonlinear setting. One notable difference compared to the main simulation is that NNR-Linear and IntelPooling slightly outperform DML-TS-NNR and DML-TS-NNR-BLM in the first two settings of this simulation.
Appendix C Additional Details for Valentine Study
Personalizing treatment delivery in mobile health is a common application for online learning algorithms. We focus here on the Valentine study, a prospective, randomized-controlled, remotely administered trial designed to evaluate an mHealth intervention to supplement cardiac rehabilitation for low- and moderate-risk patients (Jeganathan et al., 2022; Golbus et al., 2023). We aim to use smartwatch data (Apple Watch and Fitbit) obtained from the Valentine study to learn the optimal timing of notification delivery given the users’ current context.
C.1 Data from the Valentine Study
Prior to the start of the trial, baseline data was collected from each of the participants (e.g., age, gender, baseline activity level, and health information). During the study, participants are randomized to either receive a notification ($A_{t}=1$) or not ($A_{t}=0$) at each of 4 daily time points (morning, lunchtime, mid-afternoon, evening), with probability 0.25. Contextual information was collected frequently (e.g., number of messages sent in prior week, step count variability in prior week, and pre-decision point step-counts).
Since the goal of the Valentine study is to increase participants’ activity levels, we thereby define the reward, $R_{t}$, as the step count for the 60 minutes following a decision point (log-transformed to eliminate skew). Our application also uses a subset of the baseline and contextual data; this subset contains the variables with the strongest association to the reward. Table 4 shows the features available to the bandit in the Valentine study data set.
Feature Description Interaction Baseline
Phase II 1 if in Phase II, 0 o.w. $\surd$ $\surd$
Phase III 1 if in Phase II, 0 o.w. $\surd$ $\surd$
Steps in prior 30 minutes log transformed $\surd$ $\surd$
Pre-trial average daily steps log transformed $\times$ $\surd$
Device 1 if Fitbit, 0 o.w. $\times$ $\surd$
Prior week step count variability SD of the rewards in the previous week $\times$ $\surd$
Table 4: List of features available to the bandit in the Valentine study. The features available to model the action interaction (effect of sending an anti-sedentary message) and to model the baseline (reward under no action) are denoted via a $\surd$” in the corresponding column, otherwise $\times$.
For baseline variables, we use the participant’s device model ($Z_{1}$, Fitbit coded as 1), the participant’s step count variability in the prior week ($Z_{2}$), and a measure of the participant’s pre-trial activity level based on an intake survey ($Z_{3}$, with larger values corresponding to higher activity levels).
At every decision point, before selecting an action, the learner sees two state variables: the participant’s previous 30-minute step count ($S_{1}$, log-transformed) and the participant’s phase of cardiac rehabilitation ($S_{2}$, dummy coded). The cardiac rehabilitation phase is defined based on a participant’s time in the study: month 1 represents Phase I, month 2-4 represents Phase II, and month 5-6 represents Phase III.
C.2 Evaluation
The Valentine study collected the sensor-based features at 4 decision points per day for each study participant. The reward for each message was defined to be $\text{log}(0.5+x)$, where $x$ is the step count of the participant in the 60 minutes following the notification. As noted in the introduction, the baseline reward, i.e. the step count of a subject when no message is sent, not only depends on the state in a complex way but is likely dependent on a large number of time-varying observed variables. Both these characteristics (complex, time-varying baseline reward function) suggest using our proposed approach.
We generated 100 bootstrap samples and ran our contextual bandit on them, considering the binary action of whether or not to send a message at a given decision point based on the contextual variables $S_{1}$ and $S_{2}$. Each user is considered independently and with a cohesion network, for maximum personalization and independence of results. To guarantee that messages have a positive probability of being sent, we only sample the observations with notification randomization probability between $0.01$ and $0.99$. In the case of the algorithm utilizing NNR, we chose four baseline characteristics (gender, age, device, and baseline average daily steps) to establish a measure of “distance" between users. For this analysis, the value of $k$ representing the number of nearest neighbors was set to 5. To utilize bootstrap sampling, we train the Neural-Linear method’s neural network using out-of-bag samples. The neural network architecture comprises a single hidden layer with two hidden nodes. The input contains both the baseline characteristics and the contextual variables and the activation function applied here is the softplus function, defined as $\text{softplus}(x)=\log(1+\exp{(x)})$.
We performed an offline evaluation of the contextual bandit algorithms using an inverse propensity score (IPS) version of the method from Li et al. (2010), where the sequence of states, actions, and rewards in the data are used to form a near-unbiased estimate of the average expected reward achieved by each algorithm, averaging over all users.
Figure 9: (left) Unbiased estimates of the average per-trial reward for all three ablation algorithms, relative to the reward obtained under the pre-specified Valentine Study randomization policy across 20 multiple-imputed data sets. And (right) p-values from the pairwise paired t-tests.
C.3 An ablation study
We further investigate the primary contributors to the algorithm’s performance, we conducted a parallel ablation study using real-world data analysis. In this study, we compared DML-TS-NNR (as detailed in the main paper) with two additional methods for reference: DML-TS-SU and NNR-Linear. Each comparison focuses on specific elements of our algorithm to discern their impact on DML-TS-NNR’s overall performance.
The differences between these methods are illustrated in Figure. 9. In this particular dataset, NNR plays a more significant role in driving the overall superior performance of our algorithm.
C.4 Inverse Propensity Score (IPS) offline evaluation
In the implemented Valentine study, the treatment was randomized with a constant probability $p_{t}=0.25$ at each time $t$. To conduct off-policy evaluation using our proposed algorithm and the competing variations of the TS algorithm, we outline the IPS estimator for an unbiased estimate of the per-trial expected reward based on what has been studied in Li et al. (2010).
Given the logged data $\mathcal{D}=\{s_{t}=s_{t},A_{t}=a_{t},R_{t}=r_{t}\}_{t=1}^{T}$ collected under the policy $\mathbf{p}=\{p_{t}\}_{t=1}^{T}$, and the treatment policy being evaluated $\mathbf{\pi}=\{\pi_{t}\}_{t=1}^{T}$, the objective of this offline estimator is to reweight the observed reward sequence $\{R_{t}\}_{t=1}^{T}$ to assign varying importance to actions based on the propensities of both the original and new policies in selecting them.
Lemma 1 (Unbiasedness of the IPS estimator).
Assuming the positivity assumption in logging, which states that for any given $s$ and $a$, if $p_{t}(a|s)>0$, then we also have $\pi_{t}(a|s)>0$, we can obtain an unbiased per-trial expected reward using the following IPS estimator:
$\hat{R}_{\text{IPS}}=\frac{1}{T}\sum_{t=1}^{T}\frac{\pi_{t}(a_{t}|s_{t})}{p_{t}(a_{t}|s_{t})}r_{t}$ (11)
As mentioned in the previous section, we restrict our sampling to observations with notification randomization probabilities ranging from 0.01 to 0.99. This selection criterion ensures the satisfaction of the positivity assumption. The proof essentially follows from definition, we have:
Proof.
$\mathbb{E}[R_{\text{IPS}}]$ $=\mathbb{E}_{\mathbf{p}}\left[\frac{1}{T}\sum_{t=1}^{T}\frac{\pi_{t}(a_{t}|s_{t})}{p_{t}(a_{t}|s_{t})}R_{t}(a_{t},s_{t})\right]$
$=\frac{1}{T}\sum_{t=1}^{T}\frac{\pi_{t}(a_{t}|s_{t})}{p_{t}(a_{t}|s_{t})}R_{t}(a_{t},s_{t})\times p_{t}(a_{t}|s_{t})$
$=\frac{1}{T}\sum_{t=1}^{T}\pi_{t}(a_{t}|s_{t})R_{t}(a_{t},s_{t})$
$=\mathbb{E}_{\mathbf{\pi}}\left[\frac{1}{T}\sum_{t=1}^{T}R_{t}(a_{t},s_{t})\right]$
To address the instability issue caused by re-weighting in some cases, we use a Self-Normalized Inverse Propensity Score (SNIPS) estimator. This estimator scales the results by the empirical mean of the importance weights, and still maintains the property of unbiasedness.
$\hat{R}_{\text{SNIPS}}=\frac{\hat{R}_{\text{IPS}}}{\frac{1}{T}\sum_{t=1}^{T}\frac{\pi_{t}(a_{t}|s_{t})}{p_{t}(a_{t}|s_{t})}}=\frac{\sum_{t=1}^{T}\frac{\pi_{t}(a_{t}|s_{t})}{p_{t}(a_{t}|s_{t})}r_{t}}{\sum_{t=1}^{T}\frac{\pi_{t}(a_{t}|s_{t})}{p_{t}(a_{t}|s_{t})}}$ (12)
Appendix D Additional Details for the Intern Health Study (IHS)
To further enhance the competitive performance of our proposed DML-TS-NNR algorithm, we conducted an additional comparative analysis using a real-world dataset from the Intern Health Study (IHS) (NeCamp et al., 2020). This micro-randomized trial investigated the use of mHealth interventions aimed at improving the behavior and mental health of individuals in stressful work environments. The estimates obtained represent the improvement in average reward relative to the original constant randomization, averaging across stages (K = 30) and participants (N = 1553). The available IHS data consist of 20 multiple-imputed data sets. We apply the algorithms to each imputed data set and perform a comparative analysis of the competing algorithms. The results presented in Figure 10 shows our proposed DML-TS-NNR algorithm achieved significantly higher rewards than the other three competing ones and demonstrated comparable performance to the AC algorithm. These findings further support the advantages of our proposed algorithm.
D.1 Data from the IHS
Prior to the start of the trial, baseline data was collected on each of the participants (e.g., institution, specialty, gender, baseline activity level, and health information). During the study, participants are randomized to either receive a notification ($A_{t}=1$) or not ($A_{t}=0$) every day, with probability $3/8$. Contextual information was collected frequently (e.g., step count in prior five days, and current day in study).
We hereby define the reward, $R_{t}$, as the step count on the following day (cubic root). Our application also uses a subset of the baseline and contextual data; this subset contains the variables with the strongest association to the reward. Table 5 shows the features available to the bandit in the IHS data set.
Feature Description Interaction Baseline
Day in study an integer from $1$ to $30$ $\surd$ $\surd$
Average daily steps in prior five days cubic root $\surd$ $\surd$
Average daily sleep in prior five days cubic root $\times$ $\surd$
Average daily mood in prior five days a Likert scale from $1-10$$\times$ $\surd$
Pre-intern average daily steps cubic root $\times$ $\surd$
Pre-intern average daily sleep cubic root $\times$ $\surd$
Pre-intern average daily mood a Likert scale from $1-10$$\times$ $\surd$
Sex Gender $\times$ $\surd$
Week category The theme of messages in a specific week (mood, sleep, activity, or none) $\times$ $\surd$
PHQ score PHQ total score $\times$ $\surd$
Early family environment higher score indicates higher level of adverse experience $\times$ $\surd$
Personal history of depression $\times$ $\surd$
Neuroticism (Emotional experience) higher score indicates higher level of neuroticism $\times$ $\surd$
Table 5: List of features available to the bandit in the IHS. The features available to model the action interaction (effect of sending a mobile prompt) and to model the baseline (reward under no action) are denoted via a $\surd$” in the corresponding column, otherwise $\times$.
At every decision point, before selecting an action, the learner sees two state variables: the participant’s previous 5-day average daily step count ($S_{1}$, cubic root) and the participant’s day in the study ($S_{2}$, an integer from $1$ to $30$).
D.2 Evaluation
We run our contextual bandit on the IHS data, considering the binary action of whether or not to send a message at a given decision point based on the contextual variables $S_{1}$ and $S_{2}$. Each user is considered independently and with a cohesion network, for maximum personalization and independence of results. To guarantee that messages have a positive probability of being sent, we only sample the observations with notification randomization probability between $0.01$ and $0.99$. For the algorithm employing NNR, we defined participants in the same institution as their own “neighbors". This definition enables the flexibility for the value of $k$, representing the number of nearest neighbors, to vary for each participant based on their specific institutional context. Furthermore, in our study setting, we assume that individuals from the same "institution" enter the study simultaneously as a group. Due to the limited access to prior data, we are unable to build the neural linear models as in the Valentine Study.
We utilized 20 multiple-imputed data sets and performed an offline evaluation of the contextual bandit algorithms on each data set. The result is presented below in Figure 10. Similar to Section C.3, here we also compared DML-TS-NNR (as detailed in the main paper) with DML-TS-SU and NNR-Linear. The differences between these methods are illustrated in Figure. 11. In this specific dataset, both NNR and DML exhibit comparable performance individually, but their combined effect significantly enhances the overall performance of the algorithm.
Figure 10: (left) Unbiased estimates of the average per-trial reward for all four competing algorithms, relative to the reward obtained under the pre-specified Intern Health Study randomization policy across 20 multiple-imputed data sets. And (right) p-values from the pairwise paired t-tests. The dark shade in the last column indicates that the proposed DML-TS-NNR algorithm achieved significantly higher rewards than the other three competing algorithms while demonstrating comparable performance to the AC algorithm.
Figure 11: (left) Unbiased estimates of the average per-trial reward for all three ablation algorithms, relative to the reward obtained under the pre-specified Intern Health Study randomization policy across 20 multiple-imputed data sets. And (right) p-values from the pairwise paired t-tests.
Appendix E Additional Details for Algorithm 1
This appendix briefly discusses two details of Algorithm 1. The first is efficient computation of $V^{-1}$. Because $V$ is a fairly large matrix, a full matrix inversion is expensive. Fortunately, however, we can dramatically reduce the necessary computational requirements because each additional $(i,t)$ involves a rank-one perturbation to $V$. Consequently, we can apply the Sherman–Morrison formula to speed up computations. We leveraged this trick in our implementation of DML-TS-NNR for the simulation study by using efficient rank-one updates available in the SuiteSparse library (Davis & Hu, 2011).
The second detail is the requirements for the distribution, $\mathcal{D}^{TS}$:
Definition 1.
$\mathcal{D}^{TS}$ is a multivariate distribution on $\mathbb{R}^{d}$ absolutely continuous with respect to Lebesgue measure which satisfies: 1. (anti-concentration) that there exists a strictly positive probability $p$ such that for any $u\in\mathbb{R}^{d}$ with $\|u\|=1$, $\mathbb{P}(u^{\top}\eta\geq 1)\geq p$; and 2. (concentration) there exists $c$, $c^{\prime}$ positive constants such that $\forall\delta\in(0,1)$, $P(\|\eta\|\leq\sqrt{cd\log(c^{\prime}d/\delta))}\geq 1-\delta$.
While a Gaussian prior satisfies Definition 1, this approach allows us to move beyond Bayesian posteriors to generic randomized policies. As discussed in the main paper, we recommend setting $\mathcal{D}^{TS}$ to a multivariate Gaussian distribution or multivariate t-distribution. These choices simplify computations because the probability computation in (10) simplifies to an evaluation of the CDF of either (a) a univariate Gaussian or (b) a univariate t-distribution. The mean and variance of the corresponding univariate distribution can easily be worked out using the moments of a linear combination of random variables.
Appendix F Regret Bound
F.1 Double Robustness of Pseudo-Reward
Lemma 2.
If either $p_{i,t}=\pi_{i,t}$ or $f_{i,t}=r_{i,t}$, then
$\mathbb{E}\left[\tilde{R}_{i,t}^{f}|s,\bar{a}\right]$ $=\Delta_{i,t}(s,\bar{a}).$
That is, the pseudo-reward is an unbiased estimator of the true differential reward.
Proof.
Recall that
$\tilde{R}_{i,t}^{f}$ $=\frac{R_{it}-f_{i,t}(s,A_{i,t})}{\delta_{A_{i,t}=\bar{a}}-\pi_{i,t}(0|s)}+\Delta^{f}_{i,t}(s,\bar{a})$
Case I: $\pi$’s are correctly specified
Then
$\mathbb{E}\left[\frac{R_{it}}{\delta_{A_{i,t}=\bar{a}}-\pi_{i,t}(0|s)}\biggr{|}s,\bar{a}\right]$ $=r_{i,t}(s,\bar{a})-r_{i,t}(s,0)$
$=\Delta_{i,t}(s,\bar{a})$
and
$\mathbb{E}\left[\frac{f_{i,t}(s,A_{i,t})}{\delta_{A_{i,t}=\bar{a}}-\pi_{i,t}(0|s)}\biggr{|}s,\bar{a}\right]$ $=f_{i,t}(s,\bar{a})-f_{i,t}(s,0)$
$=\Delta^{f}_{i,t}(s,\bar{a})$
so that
$\mathbb{E}\left[\frac{R_{it}-f_{i,t}(s,A_{i,t})}{\delta_{A_{i,t}=\bar{a}}-\pi_{i,t}(0|s)}+\Delta^{f}_{i,t}(s,\bar{a})\biggr{|}s,\bar{a}\right]$ $=\Delta_{i,t}(s,\bar{a})-\Delta^{f}_{i,t}(s,\bar{a})+\Delta_{i,t}^{f}(s,\bar{a})$
$=\Delta_{i,t}(s,\bar{a})$
Case II: $f$ correctly specified
$\mathbb{E}\left[\frac{R_{it}}{\delta_{A_{i,t}=\bar{a}}-\pi_{i,t}(0|s)}\biggr{|}s,\bar{a}\right]$ $=\frac{1-p_{i,t}(0|s)}{1-\pi_{i,t}(0|s)}r_{i,t}(s,\bar{a})-\frac{p_{i,t}(0|s)}{\pi_{i,t}(0|s)}r_{i,t}(s,0)$
and
$\mathbb{E}\left[\frac{f_{i,t}(s,A_{i,t})}{\delta_{A_{i,t}=\bar{a}}-\pi_{i,t}(0|s)}\biggr{|}s,\bar{a}\right]$ $=\frac{1-p_{i,t}(0|s)}{1-\pi_{i,t}(0|s)}f_{i,t}(s,\bar{a})-\frac{p_{i,t}(0|s)}{\pi_{i,t}(0|s)}f_{i,t}(s,\bar{0})$
$=\frac{1-p_{i,t}(0|s)}{1-\pi_{i,t}(0|s)}r_{i,t}(s,\bar{a})-\frac{p_{i,t}(0|s)}{\pi_{i,t}(0|s)}r_{i,t}(s,\bar{0})$
and
$\mathbb{E}\left[\Delta^{f}_{i,t}(s,\bar{a})\biggr{|}s,\bar{a}\right]$ $=\Delta_{i,t}(s,\bar{a})$
$\mathbb{E}\left[\frac{R_{it}-f_{i,t}(s,\bar{A}_{i,t})}{\delta_{A_{i,t}=\bar{a}}-\pi_{i,t}(0|s)}+\Delta^{f}_{i,t}(s,\bar{a})\biggr{|}s,\bar{a}\right]$ $=\Delta_{i,t}(s,\bar{a})$
F.2 Preliminaries
Lemma 3.
Let $X$ be a mean-zero sub-Gaussian random variable with variance factor $v^{2}$ and $Y$ be a bounded random variable such that $|Y|\leq B$ for some $0\leq B<\infty$. Then $XY$ is sub-Gaussian with variance factor $v^{2}B^{2}$.
Proof.
Recall that $X$ being mean-zero sub-Gaussian means that
$P(|X|\geq t)$ $\leq 2\exp\left(-\frac{t^{2}}{2v^{2}}\right).$
Now note that
$|XY|$ $\leq|X|B$
so that if $|XY|>t$, then $|X|B>t$. Thus by monotonicity
$P(|XY|\geq t)$ $\leq P(|X|B\geq t)$
$=P\left(|X|\geq\frac{t}{B}\right)$
$\leq 2\exp\left(-\frac{t^{2}}{2B^{2}v^{2}}\right)$
as desired. ∎
Lemma 4.
If $X,Y$ are sub-Gaussian with variance factors $v_{x}^{2}$, $v_{y}^{2}$, respectively, then $\alpha X+\beta Y$ is sub-Gaussian with variance factor $\alpha^{2}v_{x}^{2}+\beta^{2}v_{y}^{2}$$\forall\alpha,\beta\in\mathbb{R}$.
Proof.
Recall the equivalent definition of sub-Gaussianity that $X,Y$ are sub-Gaussian iff for some $b,a>0$ and all $\lambda>0$
$\mathbb{E}\exp\left(\lambda(X-\mathbb{E}X)\right)$ $\leq\exp(\lambda^{2}v_{x}^{2}/2)$
$\mathbb{E}\exp\left(\lambda(Y-\mathbb{E}Y)\right)$ $\leq\exp(\lambda^{2}v_{y}^{2}/2)$
Then
$\mathbb{E}\exp\left(\lambda(\alpha X+\beta Y-\alpha\mathbb{E}X-\beta\mathbb{E}Y)\right)$ $\leq\sqrt{\mathbb{E}\exp\left(2\alpha\lambda(X-\mathbb{E}X)\right)}\sqrt{\mathbb{E}\exp\left(2\beta\lambda(Y-\mathbb{E}Y)\right)}$
$\leq\sqrt{\exp(2\alpha^{2}\lambda^{2}v_{x}^{2})}\sqrt{\exp(2\beta^{2}\lambda^{2}v_{y}^{2})}$
$=\exp((\alpha^{2}v_{x}^{2}+\beta^{2}v_{y}^{2})\lambda^{2})$
The following Lemma gives the sub-Gaussianity and variance of the difference between the pseudo-reward and its expectation. We see that in the variance, all terms except those involving the inverse propensity weighted noise variance vanish as $f_{i,t}$ becomes a better estimate of $r_{i,t}$. Note that means and variances may be implicitly conditioned on the history.
Lemma 5.
If $\pi_{i,t}$ is correctly specified and $\tilde{\sigma}_{i,t}^{2}\geq\frac{1}{c}$, the difference between the pseudo-reward and its expectation (taken wrt the action and noise) is mean zero sub-Gaussian with variance
$\textrm{Var}\left(\tilde{R}_{i,t}^{f}(s,\bar{a})\right)$ $=\frac{(r_{i,t}(s,\bar{a})-f_{i,t}(s,\bar{a}))^{2}\pi_{i,t}(0|s)+\textrm{Var}(\epsilon_{i,t})}{1-\pi_{i,t}(0|s)}$
$\qquad+\frac{(r_{i,t}(s,0)-f_{i,t}(s,0))^{2}[1-\pi_{i,t}(0|s)]+\textrm{Var}(\epsilon_{i,t})}{\pi_{i,t}(0|s)}$
$\qquad+2(r_{i,t}(s,\bar{a})-f_{i,t}(s,\bar{a}))(r_{i,t}(s,0)-f_{i,t}(s,0))$
Proof.
We need to show that it is sub-Gaussian and upper bound its variance. We write the difference as
$\tilde{R}_{i,t}^{f}(s,\bar{a})-\mathbb{E}[\tilde{R}_{i,t}^{f}|s,\bar{a}]$ $=\tilde{R}_{i,t}^{f}(s,\bar{a})-\Delta_{i,t}(s,\bar{a})$
$=\frac{R_{it}-f_{i,t}(s,A_{i,t})}{\delta_{A_{i,t}=\bar{a}}-\pi_{i,t}(0|s)}+\Delta^{f}_{i,t}(s,\bar{a})-\Delta_{i,t}(s,\bar{a})$
$=\frac{r_{i,t}(s,A_{i,t})-f_{i,t}(s,A_{i,t})+\epsilon_{i,t}}{\delta_{A_{i,t}=\bar{a}}-\pi_{i,t}(0|s)}+\Delta^{f}_{i,t}(s,\bar{a})-\Delta_{i,t}(s,\bar{a})$
Note that $|r_{i,t}(s,A_{i,t})|\leq\max\left(|r_{i,t}(s,\bar{a})|,|r_{i,t}(s,0)|\right)$ and $|f_{i,t}(s,A_{i,t})|\leq\max\left(|f_{i,t}(s,\bar{a})|,|f_{i,t}(s,0)|\right)$. Thus since $\left|\frac{1}{\delta_{A_{i,t}=\bar{a}}-\pi_{i,t}(0|s)}\right|$ is upper bounded by $c>0$, we have that $\frac{r_{i,t}(s,A_{i,t})-f_{i,t}(s,A_{i,t})}{\delta_{A_{i,t}=\bar{a}}-\pi_{i,t}(0|s)}$ is bounded and thus (not necessarily mean zero) sub-Gaussian. Since $\epsilon_{i,t}$ is sub-Gaussian, its denominator is bounded, and the remaining terms are deterministic, the entire difference between the pseudo-reward and its mean is sub-Gaussian. Now
$\textrm{Var}(\tilde{R}_{i,t}^{f}(s,\bar{a})-\mathbb{E}[\tilde{R}_{i,t}^{f}|s,\bar{a}])$ $=\textrm{Var}\left(\tilde{R}_{i,t}^{f}(s,\bar{a})\right)$
$=\mathbb{E}\left[\tilde{R}_{i,t}^{f}(s,\bar{a})^{2}\right]-\Delta_{i,t}(s,\bar{a})^{2}$ (13)
since $\mathbb{E}[\tilde{R}_{i,t}^{f}|s,\bar{a}]$ is not random. Now we expand the first term on the rhs.
$\mathbb{E}\left[\tilde{R}_{i,t}^{f}(s,\bar{a})^{2}\right]$ $=\mathbb{E}\left[\left(\frac{r_{i,t}(s,A_{i,t})-f_{i,t}(s,A_{i,t})+\epsilon_{i,t}}{\delta_{A_{i,t}=\bar{a}}-\pi_{i,t}(0|s)}+\Delta_{i,t}^{f}(s,\bar{a})\right)^{2}\right]$
$=\mathbb{E}\left[\left(\frac{r_{i,t}(s,A_{i,t})-f_{i,t}(s,A_{i,t})+\epsilon_{i,t}}{\delta_{A_{i,t}=\bar{a}}-\pi_{i,t}(0|s)}\right)^{2}\right]$
$\qquad+2\mathbb{E}\left[\frac{r_{i,t}(s,A_{i,t})-f_{i,t}(s,A_{i,t})+\epsilon_{i,t}}{\delta_{A_{i,t}=\bar{a}}-\pi_{i,t}(0|s)}\right]\Delta_{i,t}^{f}(s,\bar{a})+\Delta_{i,t}^{f}(s,\bar{a})^{2}$
$=\mathbb{E}\left[\left(\frac{r_{i,t}(s,A_{i,t})-f_{i,t}(s,A_{i,t})+\epsilon_{i,t}}{\delta_{A_{i,t}=\bar{a}}-\pi_{i,t}(0|s)}\right)^{2}\right]$
$\qquad+2(\Delta_{i,t}(s,\bar{a})-\Delta_{i,t}^{f}(s,\bar{a}))\Delta_{i,t}^{f}(s,\bar{a})+\Delta_{i,t}^{f}(s,\bar{a})^{2}$ (14)
For the first term on the rhs of Eqn. 14,
$\mathbb{E}$ $\left[\left(\frac{r_{i,t}(s,A_{i,t})-f_{i,t}(s,A_{i,t})+\epsilon_{i,t}}{\delta_{A_{i,t}=\bar{a}}-\pi_{i,t}(0|s)}\right)^{2}\right]$
$=\mathbb{E}\left[\left(\frac{r_{i,t}(s,A_{i,t})-f_{i,t}(s,A_{i,t})}{\delta_{A_{i,t}=\bar{a}}-\pi_{i,t}(0|s)}\right)^{2}\right]+\mathbb{E}\left[\left(\frac{\epsilon_{i,t}}{\delta_{A_{i,t}=\bar{a}}-\pi_{i,t}(0|s)}\right)^{2}\right]$
$=\frac{(r_{i,t}(s,\bar{a})-f_{i,t}(s,\bar{a}))^{2}+\mathbb{E}[\epsilon_{i,t}^{2}]}{1-\pi_{i,t}(0|s)}+\frac{(r_{i,t}(s,0)-f_{i,t}(s,0))^{2}+\mathbb{E}[\epsilon_{i,t}^{2}]}{\pi_{i,t}(0|s)}$
so that plugging this into Eqn. 14, we have
$\mathbb{E}\left[\tilde{R}_{i,t}^{f}(s,\bar{a})^{2}\right]$ $=\frac{(r_{i,t}(s,\bar{a})-f_{i,t}(s,\bar{a}))^{2}+\mathbb{E}[\epsilon_{i,t}^{2}]}{1-\pi_{i,t}(0|s)}+\frac{(r_{i,t}(s,0)-f_{i,t}(s,0))^{2}+\mathbb{E}[\epsilon_{i,t}^{2}]}{\pi_{i,t}(0|s)}$
$\qquad+2(\Delta_{i,t}(s,\bar{a})-\Delta_{i,t}^{f}(s,\bar{a}))\Delta_{i,t}^{f}(s,\bar{a})+\Delta_{i,t}^{f}(s,\bar{a})^{2}$
and plugging this into Eqn. 13 we obtain the variance.
$\textrm{Var}\left(\tilde{R}_{i,t}^{f}(s,\bar{a})\right)$ $=\frac{(r_{i,t}(s,\bar{a})-f_{i,t}(s,\bar{a}))^{2}+\textrm{Var}(\epsilon_{i,t})}{1-\pi_{i,t}(0|s)}+\frac{(r_{i,t}(s,0)-f_{i,t}(s,0))^{2}+\textrm{Var}(\epsilon_{i,t})}{\pi_{i,t}(0|s)}$
$\qquad+2(\Delta_{i,t}(s,\bar{a})-\Delta_{i,t}^{f}(s,\bar{a}))\Delta_{i,t}^{f}(s,\bar{a})+\Delta_{i,t}^{f}(s,\bar{a})^{2}-\Delta_{i,t}(s,\bar{a})^{2}$ (15)
Note that
$2(\Delta_{i,t}(s,\bar{a})-\Delta_{i,t}^{f}(s,\bar{a}))\Delta_{i,t}^{f}(s,\bar{a})+\Delta_{i,t}^{f}(s,\bar{a})^{2}-\Delta_{i,t}(s,\bar{a})^{2}$
$=2(\Delta_{i,t}(s,\bar{a})-\Delta_{i,t}^{f}(s,\bar{a}))\Delta_{i,t}^{f}(s,\bar{a})-(\Delta_{i,t}(s,\bar{a})-\Delta_{i,t}^{f}(s,\bar{a}))(\Delta_{i,t}(s,\bar{a})+\Delta_{i,t}^{f}(s,\bar{a}))$
$=(\Delta_{i,t}(s,\bar{a})-\Delta_{i,t}^{f}(s,\bar{a}))(\Delta_{i,t}^{f}(s,\bar{a})-\Delta_{i,t}(s,\bar{a}))$
$=-(\Delta_{i,t}(s,\bar{a})-\Delta_{i,t}^{f}(s,\bar{a}))^{2}$
and plugging this into Eqn. 15,
$\textrm{Var}\left(\tilde{R}_{i,t}^{f}(s,\bar{a})\right)$ $=\frac{(r_{i,t}(s,\bar{a})-f_{i,t}(s,\bar{a}))^{2}+\textrm{Var}(\epsilon_{i,t})}{1-\pi_{i,t}(0|s)}+\frac{(r_{i,t}(s,0)-f_{i,t}(s,0))^{2}+\textrm{Var}(\epsilon_{i,t})}{\pi_{i,t}(0|s)}$
$\qquad-(\Delta_{i,t}(s,\bar{a})-\Delta_{i,t}^{f}(s,\bar{a}))^{2}$
as desired. Now note that
$\left(\Delta_{i,t}(s,\bar{a})-\Delta_{i,t}^{f}(s,\bar{a})\right)^{2}$ $=\left(r_{i,t}(s,\bar{a})-r_{i,t}(s,0)-(f_{i,t}(s,\bar{a})-f_{i,t}(s,0))\right)^{2}$
$=\left(r_{i,t}(s,\bar{a})-f_{i,t}(s,\bar{a})-(r_{i,t}(s,0)-f_{i,t}(s,0))\right)^{2}$
$=(r_{i,t}(s,\bar{a})-f_{i,t}(s,\bar{a}))^{2}+(r_{i,t}(s,0)-f_{i,t}(s,0))^{2}$
$\qquad+2(r_{i,t}(s,\bar{a})-f_{i,t}(s,\bar{a}))(r_{i,t}(s,0)-f_{i,t}(s,0))$
so that
$\textrm{Var}\left(\tilde{R}_{i,t}^{f}(s,\bar{a})\right)$ $=\frac{(r_{i,t}(s,\bar{a})-f_{i,t}(s,\bar{a}))^{2}\pi_{i,t}(0|s)+\textrm{Var}(\epsilon_{i,t})}{1-\pi_{i,t}(0|s)}$
$\qquad+\frac{(r_{i,t}(s,0)-f_{i,t}(s,0))^{2}[1-\pi_{i,t}(0|s)]+\textrm{Var}(\epsilon_{i,t})}{\pi_{i,t}(0|s)}$
$\qquad+2(r_{i,t}(s,\bar{a})-f_{i,t}(s,\bar{a}))(r_{i,t}(s,0)-f_{i,t}(s,0))$
Corollary 1.
$\mathbb{E}_{p(s,\bar{a})}\left[\textrm{Var}\left(\tilde{R}_{i,t}^{f}(s,\bar{a})\right)\right]$ $\leq\frac{\mathbb{E}_{p(s,\bar{a})}\left[(r_{i,t}(s,\bar{a})-f_{i,t}(s,\bar{a}))^{2}\right]\pi_{i,t}(0|s)+\textrm{Var}(\epsilon_{i,t})}{1-\pi_{i,t}(0|s)}$
$\qquad+\frac{\mathbb{E}_{p(s,\bar{a})}\left[(r_{i,t}(s,0)-f_{i,t}(s,0))^{2}\right][1-\pi_{i,t}(0|s)]+\textrm{Var}(\epsilon_{i,t})}{\pi_{i,t}(0|s)}$
$\qquad+2\sqrt{\mathbb{E}_{p(s,\bar{a})}\left[(r_{i,t}(s,\bar{a})-f_{i,t}(s,\bar{a}))^{2}\right]}\sqrt{\mathbb{E}_{p(s)}\left[(r_{i,t}(s,0)-f_{i,t}(s,0))^{2}\right]}$
Proof.
It follows immediately by taking expectations and applying Cauchy Schwartz. ∎
Corollary 2.
$(r_{i,t}(\cdot,0)-f_{i,t}(\cdot,0))^{2}=o_{P}(k^{-1/2})$ and $(r_{i,t}(\cdot,\bar{a})-f_{i,t}(\cdot,\bar{a}))^{2}=o_{P}(k^{-1/2})$
Proof.
Note that
$P\left((r_{i,t}(\cdot,0)-f_{i,t}(\cdot,0))^{2}k^{1/2}>C\right)$ $\leq\frac{\mathbb{E}_{p(\mathcal{H}_{k-1})}\mathbb{E}_{p(s)}\left[(r_{i,t}(s,0)-f_{i,t}(s,0))^{2}\right]k^{1/2}}{C}$
[…]
$=\frac{\mathbb{E}_{p(\mathcal{H}_{k-1})}o_{P}(1)}{C}$
$\leq\delta\textrm{ for sufficiently large $k$}$
so that $(r_{i,t}(\cdot,0)-f_{i,t}(\cdot,0))^{2}=o_{P}(k^{-1/2})$. In the second to last line we used that since $r_{i,t},f_{i,t}$ are bounded, $\mathbb{E}_{p(s)}\left[(r_{i,t}(s,0)-f_{i,t}(s,0))^{2}\right]$ is bounded. For bounded random variables, convergence in probability implies convergence in mean. A similar result holds when using $\bar{a}$ instead of 00. ∎
Corollary 3.
$\textrm{Var}\left(\tilde{R}_{i,t}^{f}(s,\bar{a})\right)$ $\leq\frac{o_{P}(k^{-1/2})\pi_{i,t}(0|s)+\textrm{Var}(\epsilon_{i,t})}{1-\pi_{i,t}(0|s)}+\frac{o_{P}(k^{-1/2})[1-\pi_{i,t}(0|s)]+\textrm{Var}(\epsilon_{i,t})+\textrm{Var}(\epsilon_{i,t})}{\pi_{i,t}(0|s)}$
$\qquad+2o_{P}(k^{-1/2})$
Proof.
This follows immediately from the previous two Corollaries. ∎
Corollary 4.
For all $\delta>0$, there exists $C>0$ s.t. w.p. at least $1-\delta$,
$\textrm{Var}\left(\tilde{R}_{i,t}^{f}(s,\bar{a})\right)$ $\leq\frac{Ck^{-1/2}\pi_{i,t}(0|s)+\textrm{Var}(\epsilon_{i,t})}{1-\pi_{i,t}(0|s)}+\frac{Ck^{-1/2}[1-\pi_{i,t}(0|s)]+\textrm{Var}(\epsilon_{i,t})}{\pi_{i,t}(0|s)}+2Ck^{-1/2}$
$\equiv v_{k}^{2}$
Proof.
By the definition of $o_{P}$, if $(f_{i,t}-r_{i,t})^{2}=o_{P}(k^{-1/2})$ for any $A>0,\delta>0$, there exists $K$ s.t. if $K>k$, then w.p. at least $1-\delta$,
$(f_{i,t}-r_{i,t})^{2}\leq Ak^{-1/2}.$
Then for all $k\in\mathbb{N}$, since $f,r$ are bounded
$(f_{i,t}-r_{i,t})^{2}$ $\leq 16B^{2}I(k\leq K)+I(k>K)Ak^{-1/2}.$
Now choose
$C$ $\equiv\max\left(A,16B^{2}K^{1/2}\right).$
Remark 1.
In the limit as $k\rightarrow\infty$
$\tilde{\sigma}^{2}_{i,t}\textrm{Var}\left(\tilde{R}_{i,t}^{f}(s,\bar{a})\right)$ $\leq\pi_{i,t}(0|s)\textrm{Var}(\epsilon_{i,t})+(1-\pi_{i,t}(0|s))\textrm{Var}(\epsilon_{i,t})$
$=\textrm{Var}(\epsilon_{i,t})$
In the next remark, we show what the variance would be if we did not use DML and estimate $f_{i,t}\approx r_{i,t}$, but used only the inverse propensity weighted observed reward as the pseudo-reward. This was done in Greenewald et al. (2017). In this case, there are terms dependent on the mean reward that do not vanish as the number of stages goes to infinity.
Remark 2.
If we instead used as our pseudo-reward the inverse propensity weighted observed reward
$\tilde{R}_{i,t}(s,\bar{a})$ $=\frac{R_{it}}{\delta_{A_{i,t}=\bar{a}}-\pi_{i,t}(0|s)}$
this would be unbiased with variance
$v_{k}^{2}\equiv\frac{r_{i,t}(s,\bar{a})^{2}+\textrm{Var}(\epsilon_{i,t})}{1-\pi_{i,t}(0|s)}+\frac{r_{i,t}(s,0)^{2}+\textrm{Var}(\epsilon_{i,t})}{\pi_{i,t}(0|s)}-\Delta_{i,t}(s,\bar{a})^{2}$
Proof.
The unbiasedness is clear from our proof of Lemma 2. For the variance,
$\textrm{Var}\left(\frac{R_{it}}{\delta_{A_{i,t}=\bar{a}}-\pi_{i,t}(0|s)}\right)$ $=\mathbb{E}\left[\frac{(r_{i,t}(s,\bar{a})+\epsilon_{i,t})^{2}}{1-\pi_{i,t}(0|s)}+\frac{(r_{i,t}(s,0)+\epsilon_{i,t})^{2}}{\pi_{i,t}(0|s)}\right]-\Delta_{i,t}(s,\bar{a})^{2}$
$=\frac{r_{i,t}(s,\bar{a})^{2}+\textrm{Var}(\epsilon_{i,t})}{1-\pi_{i,t}(0|s)}+\frac{r_{i,t}(s,0)^{2}+\textrm{Var}(\epsilon_{i,t})}{\pi_{i,t}(0|s)}-\Delta_{i,t}(s,\bar{a})^{2}$
$=\frac{r_{i,t}(s,\bar{a})^{2}+\textrm{Var}(\epsilon_{i,t})}{1-\pi_{i,t}(0|s)}+\frac{r_{i,t}(s,0)^{2}+\textrm{Var}(\epsilon_{i,t})}{\pi_{i,t}(0|s)}$
$\qquad-(r_{i,t}(s,\bar{a})-r_{i,t}(s,0))^{2}$
$=\frac{r_{i,t}(s,\bar{a})^{2}+\textrm{Var}(\epsilon_{i,t})}{1-\pi_{i,t}(0|s)}-r_{i,t}(s,\bar{a})^{2}+\frac{r_{i,t}(s,0)^{2}+\textrm{Var}(\epsilon_{i,t})}{\pi_{i,t}(0|s)}$
$\qquad-r_{i,t}(s,0)^{2}+2r_{i,t}(s,\bar{a})r_{i,t}(s,0)$
$=\frac{r_{i,t}(s,\bar{a})^{2}\pi_{i,t}(0,s)+\textrm{Var}(\epsilon_{i,t})}{1-\pi_{i,t}(0|s)}+\frac{r_{i,t}(s,0)^{2}[1-\pi_{i,t}(0|s)]+\textrm{Var}(\epsilon_{i,t})}{\pi_{i,t}(0|s)}$
$\qquad+2r_{i,t}(s,\bar{a})r_{i,t}(s,0)$
Here we collect some important results. We first adapt an important concentration inequality for regularized least-squares estimates. Our proof follows the same basic strategy as Abbasi-Yadkori et al. (2011), but with modifications due to (a) the use of weighted least squares (b) the use of pseudo-rewards to estimate differential rewards (c) replacing the scaled diagonal regularization with Laplacian regularization.
Lemma 6 (Adapted from Theorem 2 in Abbasi-Yadkori et al. (2011)).
Let $\hat{\theta}_{k}$ be the stage $k$ regularized least squares (RLS) estimate from Algorithm 1 and $\theta^{*}$ the ground truth. For any $\delta>0$, w.p. at least $1-\delta$ the estimates $\{\theta_{k}\}_{k=0}^{\infty}$ satisfies for any $\{x_{k}\}_{k=0}^{\infty}$,
$|x_{k}^{\top}(\hat{\theta}_{k}-\theta^{*})|$ $\leq\|x_{k}\|_{V_{k-1}^{-1}}\left(v_{k}\sqrt{2\log\left(\frac{\det(V_{k-1})^{1/2}\det(V_{0})^{-1/2}}{\delta}\right)}+\lambda_{\max}(V_{0})^{1/2}kB\right),$ (16)
where $v_{k}^{2}$ is the variance factor for the difference between the pseudo-reward and its mean at stage $k$. In particular, setting $x_{k}=V_{k-1}(\hat{\theta}_{k-1}-\theta^{*})$ implies
$\|\hat{\theta}_{k}-\theta^{*}\|_{V_{k-1}}$ $\leq v_{k}\sqrt{2\log\left(\frac{\det(V_{k-1})^{1/2}\det(V_{0})^{-1/2}}{\delta}\right)}+\lambda_{\max}(V_{0})^{1/2}kB$
holds w.p. at least $1-\delta$ for all $k\geq 1$.
Proof.
Let $m_{i,t}=\tilde{\sigma}_{i,t}\phi(\textbf{x}_{i,t})$ and $\rho_{i,t}=\tilde{\sigma}_{i,t}[\tilde{R}_{i,t}^{f}(s,\bar{a})-\mathbb{E}[\tilde{R}_{i,t}^{f}|s,\bar{a}]]$. Further let
$\xi_{k}$ $\equiv\sum_{(i,t)\in\mathcal{O}_{k}}\tilde{\sigma}_{i,t}^{2}[\tilde{R}_{i,t}^{f}(s,\bar{a})-\mathbb{E}[\tilde{R}_{i,t}^{f}|s,\bar{a}]]\phi(\textbf{x}_{i,t})$
$=\sum_{(i,t)\in\mathcal{O}_{k}}\tilde{\sigma}_{i,t}[\tilde{R}_{i,t}^{f}(s,\bar{a})-\mathbb{E}[\tilde{R}_{i,t}^{f}|s,\bar{a}]]m_{i,t}$
$=\sum_{(i,t)\in\mathcal{O}_{k}}m_{i,t}\rho_{i,t}$
Then noting that $V_{k}=\sum_{(i,t)\in\mathcal{O}_{k}}m_{i,t}m_{i,t}^{\top}+V_{0}$ and letting $W_{k}$ be the diagonal matrix of weights $\tilde{\sigma}_{i,t}^{2}$, we have
$\hat{\theta}_{k}$ $=V_{k-1}^{-1}b_{k}$
$=V_{k-1}^{-1}(\xi_{k}+\Phi_{k}^{\top}W_{k}\mathbb{E}[R_{k}^{f_{k}}|\bar{a},s])$
[…]
$=V_{k-1}^{-1}\xi_{k}+V_{k-1}^{-1}\Phi_{k}^{\top}W_{k}\Phi_{k}\theta^{*}$
$=V_{k-1}^{-1}\xi_{k}+V_{k-1}^{-1}(\Phi_{k}^{\top}W_{k}\Phi_{k}+V_{0})\theta^{*}-V_{k-1}^{-1}V_{0}\theta^{*}$
$=V_{k-1}^{-1}\xi_{k}+\theta^{*}-V_{k-1}^{-1}V_{0}\theta_{k}^{*}$
and thus
$\hat{\theta}_{k}-\theta^{*}$ $=V_{k-1}^{-1}(\xi_{k}-V_{0}\theta^{*})$
which gives
$|x_{k}^{\top}\hat{\theta}_{k}-x_{k}^{\top}\theta^{*}|$ $\leq\|x_{k}\|_{V_{k-1}^{-1}}(\|\xi_{k}\|_{V_{k-1}^{-1}}+\|V_{0}\theta*\|_{V_{k-1}^{-1}})$
Now since $\xi_{k}$ is sub-Gaussian with variance factor $v_{k}^{2}$, by Theorem 1 in Abbasi-Yadkori et al. (2011), w.p. $1-\delta$,
$\|\xi_{k}\|_{V_{k-1}^{-1}}^{2}$ $\leq 2v_{k}^{2}\log\left(\frac{\textrm{det}(V_{k-1})^{1/2}\textrm{det}(V_{0})^{-1/2}}{\delta}\right)$
Further note
$\|V_{0}\theta^{*}\|_{V_{k-1}^{-1}}^{2}$ $=\theta_{k}^{*\top}V_{0}^{\top}V_{k-1}^{-1}V_{0}\theta^{*}$
$\leq\|V_{0}^{\top}V_{k-1}^{-1}V_{0}\|_{2}\|\theta^{*}\|_{2}^{2}$
$\leq\|V_{0}\|^{2}_{2}\|V_{k-1}^{-1}\|_{2}\|\theta^{*}\|_{2}^{2}$
$\leq\lambda_{\textrm{max}}(V_{0})\|\theta^{*}\|_{2}^{2}$
$\leq\lambda_{\textrm{max}}(V_{0})k^{2}B^{2}$
and thus
$|x_{k}^{\top}\hat{\theta}_{k}-x_{k}^{\top}\theta^{*}|$ $\leq\|x_{k}\|_{V_{k-1}^{-1}}\left(v_{k}\sqrt{2\log\left(\frac{\textrm{det}(V_{k-1}^{-1})^{1/2}\textrm{det}(V_{0})^{-1/2}}{\delta}\right)}+\lambda_{\max}(V_{0})^{1/2}kB\right).$
Finally, setting $x_{k}=V_{k-1}(\hat{\theta}_{k-1}-\theta^{*})$ implies
$\|\hat{\theta}_{k-1}-\theta^{*}\|^{2}_{V_{k-1}}$ $=(\hat{\theta}_{k-1}-\theta^{*})^{\top}V_{k-1}^{\top}(\hat{\theta}_{k-1}-\theta^{*})$
$=|x_{k}^{\top}\hat{\theta}_{k}-x_{k}^{\top}\theta^{*}|$
$\leq\|x_{k}\|_{V_{k-1}^{-1}}\left(v_{k}\sqrt{2\log\left(\frac{\textrm{det}(V_{k-1}^{-1})^{1/2}\textrm{det}(V_{0})^{-1/2}}{\delta}\right)}+\lambda_{\max}(V_{0})^{1/2}kB\right)$
$=\|V_{k-1}(\hat{\theta}_{k-1}-\theta^{*})\|_{V_{k-1}^{-1}}\left(v_{k}\sqrt{2\log\left(\frac{\textrm{det}(V_{k-1}^{-1})^{1/2}\textrm{det}(V_{0})^{-1/2}}{\delta}\right)}+\lambda_{\max}(V_{0})^{1/2}kB\right)$
$=\|\hat{\theta}_{k-1}-\theta^{*}\|_{V_{k-1}}\left(v_{k}\sqrt{2\log\left(\frac{\textrm{det}(V_{k-1}^{-1})^{1/2}\textrm{det}(V_{0})^{-1/2}}{\delta}\right)}+\lambda_{\max}(V_{0})^{1/2}kB\right)$
and dividing both sides by $\|\hat{\theta}_{k-1}-\theta^{*}\|_{V_{k-1}}$ implies
$\|\hat{\theta}_{k}-\theta^{*}\|_{V_{k-1}}$ $\leq v_{k}\sqrt{2\log\left(\frac{\det(V_{k-1})^{1/2}\det(V_{0})^{-1/2}}{\delta}\right)}+\lambda_{\max}(V_{0})^{1/2}kB$
Corollary 5.
If Assumption 3, then for any $\delta>0$, there exists $C>0$ s.t. w.p. at least $1-\delta$ the estimates $\{\hat{\theta}_{k}\}_{k=0}^{\infty}$ in Algorithm 2 satisfies for any $\{x_{k}\}_{k=0}^{\infty}$,
$|x_{k}^{\top}(\hat{\theta}_{k}-\theta^{*})|$ $\leq\|x_{k}\|_{V_{k-1}^{-1}}\left(\left(\frac{C}{k^{1/2}}+\sigma^{2}c^{2}\right)\sqrt{2\log\left(\frac{\det(V_{k-1})^{1/2}\det(V_{0})^{-1/2}}{\delta/2}\right)}+\lambda_{\max}(V_{0})^{1/2}kB\right),$ (17)
In particular, setting $x_{k}=V_{k-1}(\hat{\theta}_{k-1}-\theta^{*})$ implies
$\|\hat{\theta}_{k}-\theta^{*}\|_{V_{k-1}}$ $\leq\left(\frac{C}{k^{1/2}}+\sigma^{2}c^{2}\right)\sqrt{2\log\left(\frac{\det(V_{k-1})^{1/2}\det(V_{0})^{-1/2}}{\delta/2}\right)}+\lambda_{\max}(V_{0})^{1/2}kB$
holds w.p. at least $1-\delta$ for all $k\geq 1$.
Proof.
Use Corollary 3 and Lemma 6, each with $\delta/2$. Then w.p. at least $1-\delta$ the result holds. ∎
We next state a slightly modified form of a standard result of RLS (Lemma 11 in Abbasi-Yadkori et al. (2011)) that helps to guarantee that the prediction error is cumulatively small. This bounds the sum of quadratic forms where the matrix is the inverse Gram matrix and the arguments are the feature vectors. We use such terms to construct a martingale in the regret bound so that we can bound such terms and the martingale.
Proposition 1.
Let $\lambda\geq 1$ and $\gamma\geq 1$. For any arbitrary sequence $(x_{i,t})_{(i,t)\in\mathcal{O}_{k}}$, let
$V_{k+1}\equiv\sum_{(i,t)\in\mathcal{O}_{k}}\tilde{\sigma}_{i,t}^{2}\phi(x_{i,t})\phi(x_{i,t})^{\top}+V_{0},$
be the regularized Gram matrix. Then
$\sum_{k=1}^{K}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\|\phi(x_{i,t})\|_{V_{k}^{-1}}^{2}\leq 2c\log\left(\frac{\det(V_{K+1})}{\det(V_{0})}\right).$
where $c$ is a constant such that $0<\frac{1}{c}<\tilde{\sigma}_{i,t}^{2}\forall i,t\in\mathbb{N}$. Further,
$\log\det(V_{K+1})$ $\leq 2Kd\log\left(\left[\gamma+\lambda M+\frac{K+1}{8d}\right]\right)$
Proof.
By Lemma $11$ in Abbasi-Yadkori et al. (2011), we have
$\sum_{k=1}^{K}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\tilde{\sigma}_{i,t}^{2}\|\phi(x_{i,t})\|_{V_{k}^{-1}}^{2}\leq 2\log\left(\frac{\det(V_{K+1})}{\det(V_{0})}\right).$
The lower bound on the weights implies
$\sum_{k=1}^{K}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\|\phi(x_{i,t})\|_{V_{k}^{-1}}^{2}$ $\leq 2c\log\left(\frac{\det(V_{K+1})}{\det(V_{0})}\right)$
as desired. Now,
$tr(\tilde{\sigma}_{i,t}^{2}\phi(x_{i,t})\phi(x_{i,t})^{\top})$ $=\tilde{\sigma}_{i,t}^{2}tr(\phi(x_{i,t})^{\top}\phi(x_{i,t}))$
$=\tilde{\sigma}_{i,t}^{2}\|\phi(x_{i,t})^{\top}\|^{2}$
$=3\sigma_{i,t}^{2}\|x_{i,t}\|^{2}$
$\leq\frac{3}{4}$
and $tr(L_{\otimes})=tr(L)\cdot tr(I)=KMd$. Then by Hadamard’s inequality we have
$\log\det(V_{K+1})$ $\leq 2Kd\log\left(\frac{1}{2Kd}\text{tr}\left(V_{K+1}\right)\right)$
$=2Kd\log\left(\frac{1}{2Kd}\left[\gamma 2Kd+\lambda 2KMd+\frac{3}{4}K\cdot(K+1)/2\right]\right)$
$=2Kd\log\left(\left[\gamma+\lambda M+3\frac{K+1}{16d}\right]\right)$
Finally, we state Azuma’s concentration inequality which describes concentration of super-martingales with bounded differences and is useful in controlling the regret due to the randomization of Thompson sampling.
Proposition 2 (Azuma’s concentration inequality).
If a super-martingale $(Y_{t})_{t\geq 0}$ corresponding to a filtration $\mathcal{F}_{t}$ satisfies $|Y_{t}-Y_{t-1}|<c_{t}$ some constant $c_{t}$ for all $t=1,\ldots,T$ then for any $\alpha>0$:
$P\left(Y_{T}-Y_{0}\geq\alpha\right)\leq\exp\left(-\frac{\alpha^{2}}{2\sum_{t=1}^{T}c_{t}^{2}}\right).$
F.3 Proof of Theorem 1
We first decompose the regret bound
$\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\big{[}$ $\pi_{i,t}^{\star}x(S_{i,t},a_{i,t}^{*})^{\top}\theta^{*}_{i,t}-\pi_{i,t}x(S_{i,t},A_{i,t})^{\top}\theta^{*}_{i,t}\big{]}$
$=\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\big{[}$ $(\pi_{i,t}^{\star}-\pi_{i,t})x(S_{i,t},A_{i,t})^{\top}\theta^{*}_{i,t}$
$+\pi_{i,t}^{\star}\left\{x(S_{i,t},a_{i,t}^{*})^{\top}\theta^{*}_{i,t}-x(S_{i,t},A_{i,t})^{\top}\theta^{*}_{i,t}\right\}\big{]}$
$\leq\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\big{[}$ $(\pi_{i,t}^{\star}-\pi_{i,t})x(S_{i,t},A_{i,t})^{\top}\theta^{*}_{i,t}\big{]}$
$+\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\big{[}x(S_{i,t},a_{i,t}^{*})^{\top}\theta^{*}_{i,t}-x(S_{i,t},A_{i,t})^{\top}\theta^{*}_{i,t}\big{]}$
For the first term, $\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\big{[}(\pi_{i,t}^{\star}-\pi_{i,t})x(S_{i,t},A_{i,t})^{\top}\theta^{*}_{i,t}\big{]}$, one can apply a nearly identical proof to that of Greenewald et al. (2017). For the second term we need a novel strategy, and thus focus on that term. The proof follows closely from Abeille & Lazaric (2017) with several adjustments. Assumption 6 implies that we only need to consider the unit ball $\mathcal{X}=\{\|x\|\leq 1\}$. Then the second term of the regret can be decomposed into
$\underbrace{\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\left((\phi(x_{i,t}^{\star}))^{\top}\theta^{\star}-\phi(x_{i,t})^{\top}\tilde{\theta}_{k}\right)}_{R^{TS}(K)}+\underbrace{\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\left(\phi(x_{i,t})^{\top}\tilde{\theta}_{k}-\phi(x_{i,t})^{\top}\theta^{\star}\right)}_{R^{RLS}(K)}$
where $\phi(x_{i,t}^{\star})$ is the context vector under the optimal action and $\theta^{\star}$ is the true parameter value. The first term is the regret due to the random deviations caused by sampling $\tilde{\theta}_{k}$ and whether it provides sufficient useful information about the true parameter $\theta^{\star}_{k}$. The second term is the concentration of the sampled term around the true linear model for the advantage function.
Definition 2.
We define the filtration $\mathcal{F}_{k}$ as the information accumulated up to stage $k$ before the sampling procedure, that is, $\mathcal{F}_{k}=(\mathcal{F}_{1},\sigma(x_{1},r_{2},x_{2},\ldots,x_{k-1},r_{k-1}))$, and filtration $\mathcal{F}_{k}^{x}$ as the information accumulated up to stage $k$ and including the sampled context, that is, $\mathcal{F}_{t}=(\mathcal{F}_{1},\sigma(x_{1},r_{2},x_{2},\ldots,x_{k-1},r_{k-1},x_{k}))$.
Bounding $R^{RLS}(T)$.
We decompose the second term into the variation of the point estimate and the variation of the random sample around the point estimate:
$\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\left(\phi(x_{i,t})^{\top}\tilde{\theta}_{k}-\phi(x_{i,t})^{\top}\hat{\theta}_{k}\right)+\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\left(\phi(x_{i,t})^{\top}\hat{\theta}_{k}-\phi(x_{i,t})^{\top}\theta^{\star}_{k}\right)$
The first term describes the deviation of the TS linear predictor from the RLS one, while the second term describes the deviation of the RLS linear predictor from the true linear predictor. The first term is controlled by the construction of the sampling distribution $D^{TS}$, while the second term is controlled by the RLS estimate being a minimizer of the regularized cumulative squared error in (LABEL:eq:dml_finalform). In particular, the first term will be small when the TS estimate concentrates around the RLS one, while the second will be small when the RLS estimate concentrates around the true parameter vector. The next proposition gives a lower bound on the probability that, for all stages, both the RLS parameter vector concentrates around the true parameter vector and the TS parameter vector concentrates around the RLS one.
Recall that
$\beta_{k}(\delta)=v_{k}\left[2\log\left(\frac{\det(V_{k})^{1/2}}{\det(V_{0})^{1/2}\delta/2}\right)\right]^{1/2}+B$
Proposition 3.
Let $\hat{E}_{k}$ denote the event that $\hat{\theta}_{k}$ concentrates around the true parameter for all $l\leq k$, i.e., $\hat{E}_{k}=\{\forall l\leq k,\|\hat{\theta}_{l}-\theta^{\star}_{l}\|_{V_{l}}\leq\beta_{l}(\delta^{\prime})\}$. Let $\gamma_{k}(\delta)\equiv\beta_{k}(\delta^{\prime})\sqrt{cd\log\frac{c^{\prime}d}{\delta}}$et $\tilde{E}_{k}$ denote the event that $\tilde{\theta}_{l}$ concentrates around the estimated parameter for all $l\leq k$, i.e., $\tilde{E}_{k}=\{\forall l\leq k,\|\tilde{\theta}_{l}-\hat{\theta}_{l}\|_{V_{l}}\leq\gamma_{l}(\delta^{\prime})\}$. Let $E_{k}=\hat{E}_{k}\cap\tilde{E}_{k}$. Then $P(E_{k})\geq 1-\delta/2$.
Proof.
Let $\delta^{\prime}=\delta/4K$, then Lemma 6 and a union bound give us
$P(\hat{E}_{K})$ $=P(\cap_{k=1}^{K}\{\|\hat{\theta}_{k}-\theta^{\star}\|_{V_{k}}\leq\beta_{k}(\delta^{\prime})\})$
$=1-\sum_{k=1}^{K}P(\|\hat{\theta}_{k}-\theta^{\star}\|_{V_{k}}>\beta_{k}(\delta^{\prime}))$
$=1-\sum_{k=1}^{K}\delta^{\prime}=1-\delta^{\prime}K=1-\delta/4.$
Applying the TS sampling distribution and $\tilde{\theta}_{k}=\hat{\theta}_{k}+\beta_{k}(\delta^{\prime})V_{k}^{-1/2}\eta_{k}$ where $\eta_{t}$ is drawn i.i.d. from $D^{TS}$ we have
$P\left(\|\tilde{\theta}_{k}-\hat{\theta}_{k}\|_{V_{k}}\leq\beta_{k}(\delta^{\prime})\sqrt{cd\log\left(\frac{c^{\prime}d}{\delta^{\prime}}\right)}\right)=P\left(\|\eta_{k}\|\leq\sqrt{cd\log\left(\frac{c^{\prime}d}{\delta^{\prime}}\right)}\right)\geq 1-\delta^{\prime}.$
by Definition 1. A union-bound argument yields the conclusion. ∎
We can then bound $R^{RLS}(K)$ by leveraging Lemma 6 and decomposing the error via
$R^{RLS}(K)$ $\leq\sum_{k=1}^{K}\frac{1[E_{K}]}{k}\left[\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}|\phi(x_{i,t})^{\top}(\tilde{\theta}_{k}-\hat{\theta}_{k})|\right]$
$+\sum_{k=1}^{K}\frac{1[E_{K}]}{k}\left[\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}|\phi(x_{i,t})^{\top}(\hat{\theta}_{k}-\theta^{*})|\right]$
By definition of the event $E_{K}$, we have
$|\phi(x_{i,t})^{\top}(\tilde{\theta}_{k}-\hat{\theta}_{k})|1[E_{k}]\leq\|\phi(x_{i,t})\|_{V_{k}^{-1}}\gamma_{k}(\delta^{\prime}),\quad|\phi(x_{i,t})^{\top}(\hat{\theta}_{k}-\theta^{\star})|1[E_{k}]\leq\|\phi(x_{i,t})\|_{V_{k}^{-1}}\beta_{k}(\delta^{\prime})$
so from Proposition 1, we have
$\sum_{k=1}^{K}\frac{1[E_{K}]}{k}\left[\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}|\phi(x_{i,t})^{\top}(\tilde{\theta}_{k}-\hat{\theta}_{k})|\right]$
$\leq\gamma_{K}(\delta^{\prime})\sum_{k=1}^{K}\left[\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\frac{1}{k}\|\phi(x_{i,t})\|_{V_{k}^{-1}}\right]$
$\leq\gamma_{K}(\delta^{\prime})\sqrt{\sum_{k=1}^{K}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\frac{1}{k^{2}}}\sqrt{\sum_{k=1}^{K}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\|\phi(x_{i,t})\|_{V_{k}^{-1}}^{2}}$
$\leq\gamma_{K}(\delta^{\prime})\sqrt{\sum_{k=1}^{K}\frac{1}{k}}\sqrt{\sum_{k=1}^{K}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\|\phi(x_{i,t})\|_{V_{k}^{-1}}^{2}}$
$\leq\gamma_{K}(\delta^{\prime})\sqrt{H_{K}}\sqrt{\sum_{(i,t)\in\mathcal{O}_{k}}\|\phi(x_{i,t})\|_{V_{k}^{-1}}^{2}}$
$\leq\gamma_{K}(\delta^{\prime})\sqrt{H_{K}}\sqrt{2c\log\left(\frac{\det(V_{K+1})}{\det(V_{0})}\right)}.$
Using a similar derivation for the $\beta_{k}(\delta^{\prime})$ case, we obtain
$R^{RLS}(K)$ $\leq\left(\beta_{K}(\delta^{\prime})+\gamma_{K}(\delta^{\prime})\right)\sqrt{\sum_{k=1}^{K}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\frac{1}{k^{2}}}\sqrt{2c\log\left(\frac{\det\left(V_{K+1}\right)}{\det\left(V_{0}\right)}\right)}$
$\leq\left(\beta_{K}(\delta^{\prime})+\gamma_{K}(\delta^{\prime})\right)\sqrt{\sum_{k=1}^{K}\frac{1}{k}}\sqrt{2c\log\left(\frac{\det\left(V_{K+1}\right)}{\det\left(V_{0}\right)}\right)}$
$\leq\left(\beta_{K}(\delta^{\prime})+\gamma_{K}(\delta^{\prime})\right)\sqrt{H_{K}}\sqrt{2c\left[2Kd\log\left(\gamma+\lambda M+3\frac{K+1}{16d}\right)-\log\det\left(V_{0}\right)\right]}$
with probability at least $21-\delta/$ 2 by Proposition 3, where $H_{K}$ is the harmonic number. Note that $H_{K}\sim\log(K)$ for large $K$.
Bounding $R^{TS}(T)$.
Leveraging Abeille & Lazaric (2017), Definition 1 lets us bound $R^{TS}(K)$ under the event $E_{k}$. Let $\phi(x_{i,t}^{\star})(\theta)=\arg\max_{x_{i,t}\in\cal X}\theta^{\top}\phi(x_{i,t})$. Then
$R^{TS}(K)\leq\sum_{k=1}^{K}\frac{1}{k}R_{k}^{TS}1[E_{k}]\leq\frac{4\gamma_{K}(\delta^{\prime})}{d}\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\mathbb{E}\left[\|\phi(x_{i,t}^{\star})(\tilde{\theta})\|_{V_{k}^{-1}}|\mathcal{F}_{k}\right]$ (18)
We re-write the sum in (18) as:
$\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\|\phi(x_{i,t})\|_{V_{k}^{-1}}+\underbrace{\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\left(\mathbb{E}\left[\|\phi(x_{i,t}^{\star})(\tilde{\theta})\|_{V_{k}^{-1}}|\mathcal{F}_{k}\right]-\|\phi(x_{i,t})\|_{V_{k}^{-1}}\right)}_{R_{2}^{TS}}$
The first term is bounded by Proposition 1:
$\sum_{k=1}^{K}\frac{1}{k}\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\|\phi(x_{i,t})\|_{V_{k}^{-1}}\leq\sqrt{2cH_{K}\log\left(\frac{\det(V_{K+1})}{\det(V_{0})}\right)}$
The second term is a martingale by construction and so we can apply Azuma’s inequality. Under Assumption 6, so since $V_{k}\leq\frac{1}{\lambda}I$ we have
$\mathbb{E}\left[\|\phi(x_{i,t})^{\star}(\tilde{\theta})\|_{V_{k}^{-1}}|\mathcal{F}_{t}\right]-\|\phi(x_{i,t})\|_{V_{k}^{-1}}\leq\frac{2}{\sqrt{\lambda}},\quad a.s.$
This provides the upper-bound
$R^{TS}(K)\leq\frac{4\gamma_{K}(\delta^{\prime})}{d}\left(\sqrt{\frac{8K}{\lambda}\log\left(\frac{4}{\delta}\right)}+\sqrt{4cH_{K}Kd\log\left(\gamma+\lambda M+3\frac{K+1}{16d}\right)-\log\det\left(V_{0}\right)}\right).$
Overall bound.
Putting together the two bounds under a union bound argument yields the upper bound in Theorem 1; specifically, we have
$\begin{split}&\left(\beta_{K}(\delta^{\prime})+\gamma_{K}(\delta^{\prime})\left[1+\frac{4}{d}\right]\right)\sqrt{4cH_{K}Kd\log\left(\gamma+\lambda M+3\frac{K+1}{16d}\right)-\log\det\left(V_{0}\right)}\\ +&\frac{4\gamma_{K}(\delta^{\prime})}{p}\sqrt{\frac{8K}{\lambda}\log\left(\frac{4}{\delta}\right)}\end{split}$
Appendix G Notation Guide
For convenience, we summarize below some key notation used in the main paper.
  • $a=0$: control action
  • $q$: number of non-baseline treatment arms
  • $i=1,2,\ldots$: index for individuals (later in the paper, we consider $N$ individuals)
  • $t=1,2,\ldots$: index for decision points (later in the paper, we consider $T$ decision points)
  • $S_{i,t}\in\mathcal{S}$: context vector observed for individual $i$ at decision point $t$
  • $A_{i,t}\in\{0,\ldots,K\}$: action chosen for individual $i$ at decision point $t$
  • $R_{i,t}\in\mathbb{R}$: reward observed for individual $i$ at decision point $t$
  • $r_{i,t}(s,a):=\mathbb{E}\left[R_{i,t}|S_{i,t}=s,A_{i,t}=a\right]$: conditional model for the observed reward given the state and context
  • $x(s,a)\in\mathbb{R}^{p\times 1}$: feature vector containing the state and action
  • $\theta_{i,t}\in\mathbb{R}^{p}$: vector of parameters that may depend on the individual $i$ and decision point $t$. Later $\theta_{i,t}$ is written as $\theta_{i,t}=\theta+\theta_{i}^{user}+\theta_{t}^{time}$, where $\theta_{i}^{user}$ is the individual-specific but time-invariant term and $\theta_{t}^{time}$ is a shared time-specific term
  • $\delta_{a>0}$: indicator function that takes the value $1$ if $a>0$ and 00 otherwise
  • $g_{t}(s)$: baseline reward function that is observed when individuals are randomized to receive no treatment
  • $\Delta_{i,t}(s,a):=r_{i,t}(s,a)-r_{i,t}(s,0)$: linear differential reward for any action $a>0$ and state $s$
  • $\mathcal{H}_{i,t}$: history up to decision point $t$ for individual $i$
  • $\pi_{i,t}(a|s)$: probability of action $a\in[K]$ given current context $s\in\mathcal{S}$ for a fixed (implicit) history
  • $\bar{a}_{i,t}\in[K]$: potential non-baseline arm that may be chosen if the baseline arm is not chosen
  • $f_{i,t}(s,a)$: working model for the true conditional mean $r_{i,t}(s,a)$
  • $\tilde{R}_{i,t}^{f}(s,\bar{a})$: pseudo-reward given state $S_{i,t}=s$ and potential arm $\bar{a}$, which has the same expectation as the differential reward, $\Delta_{i,t}(s_{i,t},\bar{a}_{i,t})$. This term is written as $\tilde{R}_{i,t}^{f}$ in some points in the main paper, with the state and action implied.
  • $\Delta_{i,t}^{f}(s,\bar{a})$: doubly robust estimator for the differential reward
  • $I_{m}(t)\subseteq\{1,...,t\}$: the $m$-th fold, of $M$ total folds, as assigned up to time $t$. $I^{\complement}_{m}(t)$ denotes its complement.
  • $\lambda$: parameter for ridge penalization, used in $\lambda||\theta||^{2}_{2}$
  • $\tilde{\sigma}^{2}_{i,t}=\pi_{i,t}(0|s_{i,t})\cdot(1-\pi_{i,t}(0|s_{i,t}))$: weights used in the penalized regression estimation, which are inversely proportional to $var(\tilde{R}^{f}_{i,t})$
  • $\left\{d(i,j):=||\theta_{i}-\theta_{j}||^{2}_{2}\right\}_{i\neq j}$: $L_{2}$-distances for network construction
  • $G=(V,E)$: graph with nodes $V$ and edges $E$. Each node corresponds to an individual, $V:=[N]$, and $(i,j)\in E$ for the smallest $M<<N$ distances.
  • B: incidence matrix where the element $B_{v,e}$ corresponds to the $v$-th vertex (individual) and $e$-th edge
  • $L$: Laplacian
  • $\mathcal{O}_{k}=\{(i,t):i\leq k\ \&\ t\leq k+1-i\}$: the set of observed time points across all individuals at stage $k$ of the sequential requirement setting
  • $G_{user}$: nearest neighbor graph characterizing proximity in the user domain
  • $G_{time}$: nearest neighbor graph characterizing proximity in the time domain item $\Theta_{k}=\textrm{vec}[\left(\theta,\theta_{1}^{\textrm{user}},\ldots,\theta_{k}^{\textrm{user}},\theta_{1}^{\textrm{time}},\ldots,\theta_{k}^{\textrm{time}}\right)]\in\mathbb{R}^{p(2k+1)}$: the set of all parameters, including the individual-specific time-invariant parameters and the shared time-specific parameters, at stage $k$ of the sequential recruitment setting
  • $M$, $D_{user}$, $D_{time}$: terms for bound on parameters and mean estimate.